### v3

# A Theory of “Likes”

[…]
[linecolor=black,backgroundcolor=gray!10,linewidth=1pt]mybox
A Theory of “Likes”
Jean-Michel Benkert and Armin Schmutzler Benkert: Department of Economics, University of Bern, jean-michel.benkert@unibe.ch. Schmutzler: Department of Economics, University of Zurich, and CEPR, armin.schmutzler@econ.uzh.ch. This paper was previously circulated under the title “A Theory of Recommendations”. We are grateful for very helpful discussions with Piotr Dworczak, Johannes Johnen, Igor Letina, Matt Mitchell, Marc Möller, Nick Netzer, Volker Nocke, Marcus Reisinger, Egor Starkov, Jakub Steiner and seminar participants in Bern, Fribourg, Zurich, at MaCCI (2025), SAET (2023), Swiss Society of Economics of Statistics (2024), Swiss IO Day (2024) and Verein für Socialpolitik (2024).
(May 1, 2025)
This paper investigates the value of recommendations for disseminating economic information, with a focus on frictions resulting from preference heterogeneity. We consider Bayesian expected-payoff maximizers who receive non-strategic recommendations by other consumers. We show by which channels a recommendation system influences consumer demand and welfare. Our analysis reveals how the welfare-maximizing design of a recommendation system may differ from what a profit-maximizing designer would choose to do.
Keywords: recommendations, preference heterogeneity, optimal design
JEL Classification: D02, D47, D83
1 Introduction
Recommendations play an essential role for the diffusion of economic information in diverse domains. For instance, consumers may base their choices of travel destinations, restaurants, or doctors on what they have heard from their acquaintances. In the digital economy, the supply of recommendations is increasingly mediated through simple mechanisms such as “likes,” star ratings, or thumbs up/down. These coarse signals compress personal evaluations into binary or low-dimensional feedback that guides others’ decisions.
In this paper, we ask how recommendations can function when preferences are heterogeneous and product qualities cannot be objectively ranked: is a “like” a reflection of inherent quality or merely an expression of individual taste? Consider a user of a streaming platform deciding whether to “like” a documentary. Perceived quality depends on many aspects—informativeness, emotional impact, cinematography, pacing—and different viewers weigh these aspects differently. Moreover, consumers who send recommendations often have little incentive to consider how their like will be interpreted, while the receiver may depend on it to avoid poor matches. And even a thoughtful sender cannot convey all relevant nuances through a coarse signal. Yet, recommendation systems are pervasive and people routinely rely on them. This raises fundamental questions: When are such systems useful? What shapes the value of the recommendations they generate for consumers? And how might the design incentives of platforms diverge from the interests of consumers?
The streaming example highlights three core features of recommendations that motivate our modeling approach. First, products and preferences are heterogeneous: Products come in different versions, and consumers value them differently. Second, stakes are asymmetric: Avoiding misunderstanding is more important for the receiver than for the sender. Third, recommendations are inherently coarse: The sender cannot fully convey all relevant aspects of the product. [^1] These features arise in virtually all recommendation contexts, including personal interactions with a friend. However, they are particularly pronounced on online platforms, where recommendations are typically stripped of context, reduced to simple signals, and come from (largely) anonymous senders.
Our model captures these features in a parsimonious way. An agent must decide between two products, which are two-dimensional objects differing in whether they have high or low quality in each dimension. This gives rise to four distinct product versions. There is a continuum of agent types. All types agree that a product with high quality in both dimensions is worth choosing (objectively good), and that one with low quality in both is not (objectively bad)—we refer to such products as uncontroversial. By contrast, controversial products have high quality in one dimension and low quality in the other. An agent prefers the controversial product that scores highly in the dimension he cares more about (subjectively good) to the one that scores highly in the less important dimension (subjectively bad).
We assume that products are experience goods so that, before consumption, each agent only knows the probability of each product version. However, he obtains a recommendation from a previous consumer of one of the products. The recommendation reveals whether the sender had a consumption experience that was sufficiently good (a buy recommendation), with payoffs above an exogenous threshold level $R$, or not (a don’t-buy recommendation). Therefore, the sender (she) always gives a buy recommendation for a good product and a don’t-buy recommendation for a bad product, independently of the threshold $R$. In addition, she sends a buy recommendation for a controversial product if her preferences for the product are strong enough.
This approach combines the features of recommendations discussed above in a stylized way: In line with her low stakes in the issue, the sender only provides a non-strategic and coarse description of her own consumption experience. [^2] In reality, recommendations can convey more content than whether a buyer liked a product or not, for instance, by including verbal descriptions. However, given the typical complexity of products, even relatively detailed recommendations will usually only transmit coarse information. By contrast, the receiver (he) of the recommendation, with his high stakes in the issue, will carefully evaluate the informational content of the recommendation using Bayesian inference, taking into account the distribution of product qualities, his own preferences, and those of the population. Further, our decision to focus only on buy and don’t-buy recommendations is a particularly simple way of capturing the ordinal nature of most (if not all) recommendation systems, which typically ask for giving “likes” or “dislikes”, rankings from 1 to 5 stars, and so forth. [^3] We do consider an extension to more than two types of recommendations in Section 6.
Recommendations contain objective and subjective content. Intuitively, from a buy recommendation a receiver learns that the product is not objectively bad and the chances that it is objectively good have increased. All receivers agree that this information makes it more attractive to buy the product. Moreover, the buy recommendation potentially influences the relative probabilities of controversial products, making it more likely that the product is the one that most consumers prefer. For consumers whose tastes are aligned with the majority of the population, this subjective content reinforces the positive objective content conveyed by the message. Thus, such receivers will definitely accept the recommendation. Contrarians, however, who understand that their tastes are not aligned with those of the rest of the population may decide not to accept the recommendation if the negative subjective content dominates the positive objective content.
Building on these simple insights into receiver behavior, we characterize both the value created by the recommendation system and the resulting demand for the recommended product. A recommendation system influences two groups of buyers: those who follow a buy recommendation (i.e., types who accept recommendations) and those who buy despite receiving a don’t-buy recommendation (i.e., types who reject recommendations). Each of these two groups may benefit from the recommendation system. For types whose preferences are sufficiently aligned with the general population, accepting buy recommendations provides value by successfully guiding them toward a product that outperforms the outside option in expectation. By contrast, for types whose preferences are misaligned with the general population, don’t-buy recommendations can provide a valuable positive signal about the fit of the product with their subjective preference.
We then turn to the question of optimal design. In practice, platforms may exert influence over how users give recommendations. For example, Uber employs a five-star system but informs riders that “5 stars means the ride went without issues,” while “4 stars or less” prompt requests for feedback or suggested improvements (Uber, 2024). [^4] In the case of Airbnb, a service provider notes that “A 5-star rating on Airbnb means that everything was as expected and that guests are happy with their stay,” while “4 stars is also considered a positive rating, but it’s not enough,” and that “3 stars or less are considered the negative feedback” (Hospitable, 2024). This suggest that platforms do not merely choose a recommendation format but also influence how users map experiences into ratings. In our framework, this mapping is governed by the threshold $R$: by shaping what counts as a positive signal, the platform effectively controls how selectively buy recommendations are given.
It is natural to assume that the platform’s choice of $R$ is guided by profit-maximization motives. As a result, the platform may aim to maximize demand, either because it receives fixed transaction fees or because higher traffic yields greater advertising revenue. We investigate the resulting choice of $R$ and compare it to what would be optimal from a consumer perspective.
We begin with a setting in which preference heterogeneity is purely idiosyncratic noise, so that the population is symmetric with respect to their valuations for the controversial products. In this case, there is no systematic preference bias in the population, so recommendations have no subjective content. We find that the value for consumers generated by the recommendation system is monotone in the threshold $R$. If good products are more likely than bad ones, the value increases with $R$, and the optimal threshold induces consumers to issue buy recommendations only when the product is objectively good. If bad products are more likely than good ones, the value decreases with $R$, and, from a consumer perspective, buy recommendations should be given unless the product is objectively bad. Either way, the designer optimally renders recommendations essentially objective in content.
In contrast with the value of the recommendation system for consumers, the demand for the recommended product is always decreasing in the threshold. Thus, a platform that seeks to maximize sales favors a low threshold, so that buy recommendations are issued except for objectively bad products. As a result, the objectives of maximizing value and maximizing demand are aligned in some environments but stand in sharp conflict in others.
We conclude the analysis of a symmetric population by examining how both the value of the recommendation system and the demand it induces depend on the economic environment. We find nuanced effects of changes in the prevalence of controversial products and in the degree of polarization in the population.
While the symmetric benchmark is analytically clean and provides useful intuition, it leaves no room for subjective content of recommendations. We therefore extend the analysis to asymmetric populations, where the controversial product versions are not valued equally on average. In this setting, value maximization no longer necessarily requires an extreme threshold which conveys the message whether a product is objectively good or objectively bad. Instead, it may be desirable to use interior thresholds which can effectively communicate whether a product is in line with majority tastes.
Next, we turn to demand maximization in the asymmetric setting. Using a combination of theoretical and numerical analysis, we find that a demand-maximizing platform usually prefers a low threshold. While this behavior may coincide with value maximization in some environments, the two objectives generally diverge in the presence of asymmetries.
Finally, various extensions of the model lead to further insights. First, we allow the sender and receiver populations to differ, as in cases where recommendations come from, say, early adopters. Our main results remain robust, and we discuss how the designer may want to shape the sender distribution. Second, we consider richer recommendation systems with more than two levels. While this added granularity might seem beneficial, we find that extreme thresholds remain optimal for value maximization in symmetric populations, and that additional levels do not help a demand-maximizing platform either. Third, we analyze the case where receivers observe multiple independent recommendations from a symmetric population. Perhaps surprisingly, access to more signals does not systematically increase value or demand: receiver behavior with more than one signal often mirrors that under a well-calibrated single-recommendation system.
Research on recommendations in the presence of preference heterogeneity is surprisingly scarce. An exception is Acemoglu et al. (2022), which argues that consumers with strong initial preferences for a product are more likely to purchase it so that the selection of recommenders is biased. Receivers should recognize this bias rather than accepting recommendations at face value. Contrary to Acemoglu et al. (2022), we focus on the determinants of the value and induced demand of recommendation systems and the design of those systems rather than the possibility and speed of learning.
More broadly related, various papers deal with public goods problems in the provision of recommendations, which lead to insufficient exploration and reviewing, in particular for lesser-known niche products (e.g., Kremer et al., 2014, and Che and Hörner, 2018). [^5] Expanding on these papers, Vellodi (2022) considers the impact on market structure, showing how unbiased reviews might disadvantage entrants due to the “cold start” problem, thereby reinforcing the monopoly position of incumbents. These studies also investigate how appropriate information management can mitigate these public goods problems.
While truthful recommendations are plausible in our setting, there are contexts where recommenders, firms, and platforms may have strategic incentives to bias recommendations. [^6] Bolton et al. (2013) have shown that eBay users often refrained from leaving negative feedback to avoid retaliation. Building on such concerns, Tadelis (2016) discusses how these strategic incentives shaped eBay’s platform design—for instance, the decision to disable seller evaluations of buyers—and explores broader implications for the credibility of recommendation and reputation systems. Like us, Chakraborty and Harbaugh (2010) assume that the sender possesses multidimensional information. They show how this helps her influence the behavior of the receiver in spite of a common bias. Bourreau and Gaudin (2022) consider incentives for firms to manipulate recommendations in a setting with heterogeneous products. They focus on how a platform wants to use recommendations to shift consumers to more desirable suppliers and how this affects pricing. Abstracting from product heterogeneity, Peitz and Sobolev (forthcoming) consider profit-maximizing firms that engage in sophisticated strategies to inflate their ratings. [^7] Johnen and Ng (2024) analyze a related phenomenon in a model without preference heterogeneity. They find that firms initially lower prices to attract positive ratings, later leveraging these ratings to justify price increases—effectively reducing the reliability of ratings as a quality indicator. This leads to a reduction in welfare due to inefficient product–consumer matches. The rating inflation is somewhat akin to our result that a low recommendation threshold can maximize demand and thus profits. However, in our model, this need not come at the expense of consumer welfare: when products and preferences are heterogeneous, the objectives of value and demand maximization can align, making low thresholds potentially optimal from both perspectives.
Section 2 introduces the model. In Section 3, we describe the conditions under which different receiver types accept recommendations and use this to characterize the value and demand induced by a recommendation system, respectively. Then, we in turn consider symmetric (Section 4) and asymmetric (Section 5) populations to study the optimal design of recommendation systems, both in terms of creating value and inducing demand. Section 6 discusses our results and considers extensions. Section 7 concludes. All proofs are relegated to the appendix.
2 Model
Products and Payoffs
We consider a consumer’s choice between two products. We assume that each product is fully characterized by a two-dimensional quality vector, where $Q_{d}=1$ and $Q_{d}=0$ capture high and low quality, respectively, in dimension $d\in\{1,2\}$. Accordingly, four different product versions exist. The consumer cannot observe product quality before consumption, so the alternatives are identical for him ex ante. However, he knows the prior distribution of the quality vectors, referred to as quality distribution for brevity. Consumers have heterogeneous preferences fully characterized by their type $i\in[-1/2,1/2]$. A consumer of type $i$ receives a payoff of
$v(Q_{1},Q_{2},i)=(1/2+i)Q_{1}+(1/2-i)Q_{2}$
from consuming $(Q_{1},Q_{2})$. [^8] For convenience, we identify a product version with the corresponding quality vector. Thus, possible payoffs range from 0 to 1, with all consumers obtaining the highest payoff, $1$, from consuming version $(1,1)$ and the lowest payoff, 00, from consuming $(0,0)$. Therefore, consumers agree that product versions $(1,1)$ and $(0,0)$ are (objectively) good and bad, respectively. By contrast, the payoff from consuming $(1,0)$ or $(0,1)$ depends on a consumer’s type $i$ so that the assessment of $(1,0)$ and $(0,1)$ is subjective. We thus refer to products with these quality vectors as controversial. Consumers of type $i>0$ prefer $(1,0)$ to $(0,1)$, and vice versa for $i<0$. Those of type $i=0$ are indifferent between these two versions.
The distribution of consumer types is denoted by $F$. For now, we leave the distribution unspecified but for simplicity assume continuity and full support. Finally, the quality distribution is given as $\mathbf{q}=(q_{H},q_{1},q_{2},q_{L})$ where
$\Pr[(Q_{1},Q_{2})=(1,1)]$ $=q_{H},$ $\Pr[(Q_{1},Q_{2})=(1,0)]$ $=q_{1},$
$\Pr[(Q_{1},Q_{2})=(0,0)]$ $=q_{L},$ $\Pr[(Q_{1},Q_{2})=(0,1)]$ $=q_{2}.$
We assume that $\mathbf{q}>0$ unless mentioned otherwise. We sometimes use the following self-explanatory terminology:
Definition 1.
The prevalence of controversial products is given as $Q:=\frac{q_{1}+q_{2}}{2}$. The odds of a good product are $\sigma:=q_{H}/q_{L}$.
Recommendations and Updating
We assume that one of the products comes with a recommendation. The receiver has to choose between this product and the alternative without a recommendation. We distinguish between buy recommendations ($r=B$) and don’t-buy recommendations ($r=D$). Moreover, $r=0$ corresponds to the case without recommendation. In the benchmark model, we assume that the senders (she) and receivers (he) of recommendations are both randomly drawn from $F$.
To eliminate any scope for strategic behavior of the sender, we assume that she mechanically gives a buy recommendation if she obtained a payoff of at least $R\in(0,1)$ and a don’t-buy recommendation otherwise. [^9] We exclude the threshold $R=0$ as this would render the recommendation uninformative. For symmetry, we have also excluded $R=1$, but all results would essentially go through if we allowed $R=1$. As good products (version $(1,1)$) yield a payoff of 1, they always result in a buy recommendation, whereas bad products (version $(0,0)$) always yield a don’t-buy recommendation. The sender’s type determines the recommendation for a controversial product: After having consumed $(1,0)$, a sender of type $i$ gives a buy recommendation if and only if $i\geq R-1/2$, whereas only senders with $i\leq 1/2-R$ give a buy recommendation after consuming $(0,1)$. All told, a recommendation can reflect objective and subjective considerations.
Figure 1: Illustration of Sender Behavior
1/212-1/2- 1 / 2001/4141/41 / 41/2121/21 / 2Product version $(1,0)$1/212-1/2- 1 / 21/414-1/4- 1 / 4001/2121/21 / 2Product version $(0,1)$
We assume that receivers are Bayesian, updating beliefs about product quality after having obtained a recommendation while taking into account the uncertainty about its informational content. The probability that a randomly chosen sender gives a buy or don’t-buy recommendation, respectively is given as
$\pi^{B}(R)$ $:=q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R),$
$\pi^{D}(R)$ $:=1-\pi^{B}(R)=q_{1}(1-\phi_{1}(R))+q_{2}(1-\phi_{2}(R))+q_{L}.$
Here $\phi_{1}(R):=1-F(R-1/2)$ and $\phi_{2}(R):=F(1/2-R)$ denote the probability of a buy recommendation conditional on the product version being $(1,0)$ or $(0,1)$, respectively. In words, the probability of receiving a buy recommendation from a randomly chosen sender is given by the sum of the probability that a product is objectively good and the probability that it is subjectively good enough from a randomly chosen sender’s perspective. In the appendix, we present the receiver’s posterior $\mathbf{p}^{r}(R):=(p_{H}^{r}(R),p_{1}^{r}(R),p_{2}^{r}(R),p_{L}^{r}(R))$ after a buy recommendation $r=B$ or a don’t-buy recommendation $r=D$ [^10] To ease notation, we will typically drop the dependence of these posteriors and other variables on $R$.
Finally, we use the following summary terminology:
Definition 2.
$\mathcal{E}:=(\textbf{q},F)$ defines a decision environment. A recommendation system $\mathcal{R}$ consists of a decision environment $\mathcal{E}$ and a recommendation threshold $R\in(0,1)$.
3 Preliminaries: Receiver Behavior, Value and Demand
In this section, we begin by asking under which circumstances a receiver will accept recommendations, depending on his type $i$ and the recommendation system $\mathcal{R}$. Then, we will use this result to characterize the value of a recommendation system and determine the demand induced by the recommendation system.
3.1 Receiver Behavior
Denote by $U^{B}(i)$ and $U^{D}(i)$ player $i$’s expected payoff of buying a product with a buy and don’t-buy recommendation, respectively, and by $U^{0}(i)$ the expected payoff of not buying, i.e., of buying the alternative product with no recommendation. We have
$U^{r}(i)$ $:=p_{H}^{r}+(1/2+i)p_{1}^{r}+(1/2-i)p_{2}^{r},\quad r\in\{B,D\};$
$U^{0}(i)$ $:=q_{H}+(1/2+i)q_{1}+(1/2-i)q_{2}.$
Receiver $i$ will accept a buy recommendation if $U^{B}(i)\geq U^{0}(i)$ or equivalently
$p_{H}^{B}-q_{H}+\frac{p_{1}^{B}-q_{1}}{2}+\frac{p_{2}^{B}-q_{2}}{2}\geq i[(p_{2}^{B}-q_{2})-(p_{1}^{B}-q_{1})].$ (1)
Intuitively, a buy recommendation conveys the following information to the receiver:
  1. (I)
    The probability of a bad product has fallen to zero.
  2. (II)
    The likelihood of a good product relative to the controversial products has increased.
  3. (III)
    The odds between the controversial products has changed.
Similarly, the receiver will accept a don’t-buy recommendation if $U^{0}(i)\geq U^{D}(i)$, and the effects of these recommendations are analogous. For more compact formulations, we introduce the following terminology.
Definition 3.
  1. (i)
    The objective effect of a recommendation $r$ is defined as
    $\Delta_{O}^{r}:=p_{H}^{r}-q_{H}+\frac{p_{1}^{r}-q_{1}}{2}+\frac{p_{2}^{r}-q_{2}}{2}.$ (2)
  2. (ii)
    The subjective effect of a recommendation is defined as
    $\Delta_{S}^{r}:=(p_{2}^{r}-q_{2})-(p_{1}^{r}-q_{1}).$ (3)
We can now restate condition (1) about accepting a buy recommendation as
$\Delta_{O}^{B}\geq\Delta_{S}^{B}i.$ (4)
[…]
For the unbiased receiver $i=0$, equation (4) simplifies to $\Delta_{O}^{B}\geq 0$. As this receiver does not care about the shift in probabilities between the two controversial versions of the products, $\Delta_{O}^{B}\geq 0$ reflects the two objective effects (I) and (II). For $i\neq 0$, the interpretation of $\Delta_{O}^{B}\geq 0$ is more nuanced. If $q_{1}=q_{2}$, the expected payoff of a controversial product is type-independent, ensuring that $\Delta_{O}^{B}$ fully captures the objective benefits of a buy recommendation. If $q_{1}\neq q_{2}$, the size of these benefits depends on the receiver’s type. [^12] Intuitively, if one controversial version (e.g., $(1,0)$) is ex ante more likely, learning that a product is controversial rather than bad is more valuable for those who prefer $(1,0)$ over $(0,1)$, and vice versa for the opposite preference. Similarly, learning that a product is good rather than controversial is more valuable for those who favor the less likely version. Nonetheless, the sign of $\Delta_{O}^{B}$ remains a reliable indicator of the objective effect, as all types agree that (I) and (II) are beneficial. We can now state our first result.
Proposition 1.
For a given recommendation system $\mathcal{R}$, a receiver of type $i$
  1. (i)
    accepts buy recommendations if and only if he accepts don’t-buy recommendations;
  2. (ii)
    accepts (buy and don’t-buy) recommendations if and only if $\Delta_{O}^{B}\geq i\Delta_{S}^{B}$.
Intuitively, a receiver accepts recommendations if he thinks that they have sufficient objective content and/or that his preferences are sufficiently aligned with the sender’s. Conversely, a receiver does not accept recommendations if he thinks that they do not have sufficient objective content and/or that his preferences are not sufficiently aligned with the sender’s. Whether the recommendation is positive or negative does not matter for this assessment. Denoting the type who is indifferent between accepting and not accepting recommendations as
$\tilde{i}:=$ $\frac{\Delta_{O}^{B}}{\Delta_{S}^{B}}\text{ if }\Delta_{S}^{B}\neq 0\text{,}$
we can summarize optimal receiver behavior as follows.
Corollary 1.
  1. (i)
    If $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, then all $i\in[-1/2,1/2]$ accept recommendations.
  2. (ii)
    If $\Delta_{S}^{B}<-2\Delta_{O}^{B}$, then all $i\geq\tilde{i}$ accept recommendations.
  3. (iii)
    If $\Delta_{S}^{B}>2\Delta_{O}^{B}$, then all $i\leq\tilde{i}$ accept recommendations.
The intuition for Corollary 1, the proof of which is omitted, is straightforward. As long as the subjective content of the recommendation is small (case (i)), even the most extreme receivers satisfy the cut-off condition of Proposition 1(ii). As the subjective content increases, consumers who are not sufficiently aligned with the population at large will reject recommendations (cases (ii) and (iii)).
The interplay between objective and subjective effect leads to a stark result for extreme recommendation thresholds:
Proposition 2.
Suppose $R\rightarrow 1$ or $R\rightarrow 0$. Then, all types accept recommendations irrespective of the product distribution.
In the limits, recommendations contain only objective information. As $R\to 1$, a buy recommendation guarantees the product is good, while a don’t-buy recommendation implies it is not. Since no information is revealed about the relative likelihood of controversial products, all types value the recommendation equally and accept it.
Next, we discuss how the analysis would become trivial in the absence of either controversial products or preference heterogeneity. If there were no controversial products ($q_{1}=q_{2}=0$), recommendations would simply reveal whether the product is good or bad. In this case, recommendations would contain only objective information, eliminating any friction in the system. If there was no preference heterogeneity, the sender’s recommendations would convey objective information about whether the product is good, bad, or controversial. Without heterogeneity in receiver preferences, the subjective effect of a recommendation—how it shifts beliefs between different controversial product versions—would be irrelevant, as all receivers would interpret the information in the same way. As a result, the sender could fully align recommendations with the receivers’ interests by adjusting the recommendation threshold, effectively eliminating any tension. Hence, it is the interplay between controversial products or preference heterogeneity that makes the design of the recommendation system nontrivial and interesting.
3.2 Value of the Recommendation System
Having determined receiver behavior, we can now characterize the value of the recommendation system. Without a recommendation system, all products are identical ex ante, with an expected payoff based on the prior distribution of product qualities. Hence, the expected payoff of a receiver $i$ is
$V^{0}(i):=U^{0}(i).$ (5)
Now suppose one of the products comes with a buy or don’t-buy recommendation, so the receiver has to update his belief about the quality distribution. A receiver who accepts recommendations chooses this product (rather than one of the alternatives) only if it comes with a buy recommendation. With a recommendation system, the expected payoff of a receiver $i$ who accepts recommendations is thus
$V^{A}(i):=\pi^{B}U^{B}(i)+(1-\pi^{B})U^{0}(i).$ (6)
The first term on the right-hand side comes from buy recommendations, the second one from don’t-buy recommendations.
A receiver who does not accept recommendations will buy the recommended product if it comes with a don’t-buy recommendation. Otherwise, he will buy one of the other products. Hence, the expected payoff of such a receiver is
$V^{N}(i):=$ $\pi^{B}U^{0}(i)+(1-\pi^{B})U^{D}(i).$ (7)
A buy recommendation does not create value for receivers who do not accept it, as such types choose the alternative for which no recommendation is available. Instead, the don’t-buy recommendation, which is given with probability $(1-\pi^{B})$, does influence the behavior of the types who do not accept it: Rather than buying an alternative based on prior probabilities, updating leads them to choose the product for which they received the explicit recommendation not to buy.
We now define the value of a recommendation system in terms of the expected welfare increase it creates for receivers. [^13] This definition is in line with the definition of a valuable signal in Kamenica and Gentzkow (2011).
Definition 4.
$V(\mathcal{R})$, the value of a recommendation system $\mathcal{R}$, is the expected increase in payoffs resulting from the existence of recommendations, where the expectation is taken over all pairs of independently drawn senders and receivers.
Using Definition 4, we can characterize $V(\mathcal{R})$ as follows:
vertical space
[…]
vertical space
Importantly, the definition captures the idea that the increase in expected utility is measured before drawing the sender-receiver pair. Part (i) relies on Corollary 1, which implies that all receivers accept recommendations when $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, as the objective effect dominates. Hence, by equation (6), the recommendation system creates value only in case of a buy recommendation, as a don’t-buy recommendation leads to the same expected payoff as without a recommendation system. Parts (ii) and (iii) refer to cases where the subjective content of the messages matters more. Then, not all receivers accept the recommendation, and the system may create value for both receivers who accept recommendations as well as those who do not. We restate the value $V(\mathcal{R})$ to better understand how it reflects the role of the objective and subjective content of recommendations:
Proposition 3.
The value $V(R)$ of the recommendation system $\mathcal{R}$ is: [^14] As we focus on $R$ as the design parameter, we often write $V(R)$ instead of $V(\mathcal{R})$.
vertical space
(i) $\pi^{B}[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]]$ if $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$; (ii) $(1-\pi^{B})F(\tilde{i})\left[\Delta_{O}^{D}-\Delta_{S}^{D}\mathbb{E}[i\mid i\leq\tilde{i}]\right]+\pi^{B}(1-F(\tilde{i}))\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i\mid i\geq\tilde{i}]\right]$ if $\Delta_{S}^{B}<-2\Delta_{O}^{B}$; (iii) $\pi^{B}F(\tilde{i})\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i\mid i\leq\tilde{i}]\right]+(1-\pi^{B})(1-F(\tilde{i}))\left[\Delta_{O}^{D}-\Delta_{S}^{D}\mathbb{E}[i\mid i\geq\tilde{i}]\right]$ if $\Delta_{S}^{B}>2\Delta_{O}^{B}$.
The positive payoff effect of the recommendation in case (i) reflects the type-independent objective part $\Delta_{O}^{B}$ and the subjective part $\Delta_{S}^{B}$, which is weighted by the average type. As the subjective component $\Delta_{S}^{B}$ becomes stronger in absolute terms, not all receiver types will follow the recommendation anymore, and the analysis becomes more nuanced. To illustrate this, consider case (iii) where $\Delta_{S}^{B}>2\Delta_{O}^{B}$. In this case, a buy recommendation shifts beliefs toward the controversial product version that is less preferred by receivers $i>0$, which implies that they may decide not to follow the recommendation. Only those whose preferences are sufficiently aligned with the sender’s—specifically, those with types below a certain threshold $\tilde{i}$—still choose to buy. These types represent a fraction $F(\tilde{i})$ of the population. For them, the recommendation carries both objective value (because the product is likely better than the outside option) and subjective value (because the controversial version it signals is relatively well aligned with their preferences). This subjective value is captured by the average type among those who follow the recommendation, $\mathbb{E}[i\mid i\leq\tilde{i}]$.
Even those types who reject the buy recommendation may benefit. Types above $\tilde{i}$, who choose not to follow the recommendation, infer that the product is likely of the version they personally dislike—yet this very inference tells them something useful. When these same types receive a don’t-buy recommendation, they update in favor of the controversial product version they prefer, and may end up buying despite, or, rather, because of the negative signal. So, while the objective signal says “don’t buy”, the subjective shift in beliefs may outweigh it for some types. As a result, even rejected recommendations can create value, precisely because they help those whose preferences diverge from the majority make more informed choices.
3.3 Demand for the Recommended Product
A platform does not necessarily design a recommendation system to maximize the value it creates for consumers. Rather, the objective may be to maximize profits, which in some circumstances amounts to maximizing demand. [^15] For instance, assuming that a platform has committed to a fixed transaction fee, the profit-maximizing platform will want to to adjust its recommendation system to maximize demand. Whether a receiver buys the recommended product depends on the content of the recommendation and whether the receiver accepts or rejects the recommendation. Therefore, demand may comprise receivers who accept and receivers who reject recommendations. Using Corollary 1, the demand for the recommended product $D(R)$ can be expressed analogously to the value of the recommendation system in Proposition 3 as
$D(R)=\begin{cases}\pi_{B}(R)&\text{if }|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}\\ \pi_{B}(R)F(\tilde{i}(R))+(1-\pi_{B}(R))(1-F(\tilde{i}(R)))&\text{if }\Delta_{S}^{B}>2\Delta_{O}^{B},\\ \pi_{B}(R)(1-F(\tilde{i}(R)))+(1-\pi_{B}(R))F(\tilde{i}(R))&\text{if }\Delta_{S}^{B}<-2\Delta_{O}^{B},\end{cases}$ (8)
where we have made explicit that the indifferent receiver $\tilde{i}$ and the probability of a buy recommendation $\pi_{B}$ are functions of the threshold $R$. As for the value above, the relation between $\Delta_{S}^{B}$ and $\Delta_{O}^{B}$ determines which types accept a recommendation. In the first case, all types accept a recommendation so that demand is given by the probability of a buy recommendation $\pi_{B}(R)$. In the other cases, demand comprises both types that accept a recommendation and types that do not. Importantly, types that do not accept recommendations buy the recommended product following a don’t-buy recommendation, thus contributing to demand.
4 Symmetric Populations
In this section, we consider populations that are symmetrically distributed. Hence, there is no systematic preference bias for either of the controversial products in the population, so the preference heterogeneity reflects pure noise. We will first consider value-maximizing recommendation systems. Then, we turn to demand-maximizing recommendation systems. Throughout this section, we invoke the following symmetry assumption on the preference distribution.
Assumption 1.
$F(-i)=1-F(i)$ for all $i\in[-1/2,1/2]$.
4.1 Value-Maximizing Recommendation Systems
Under Assumption 1, the value of the recommendation system can be expressed particularly concisely, as evidenced by the following result:
Proposition 4.
If Assumption 1 holds, all types accept recommendations and
$V(R)=\pi^{B}\Delta_{O}^{B}\text{.}$
Intuitively, given the symmetry assumption, both controversial product versions lead to a buy recommendation with the same probability. Hence, a recommendation does not shift probabilities between the controversial products. As the other two (objective) effects of a buy recommendation on payoffs are positive for all types, the buy recommendation increases expected payoffs for any type, so that all types accept recommendations. Proposition 3 thus implies $V(R)=\pi^{B}\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]\right]$. Further, because of the symmetry of the population, $\mathbb{E}[i]=0$, so $V(R)=\pi^{B}\Delta_{O}^{B}$ as required.
We now use our characterization of the value of a recommendation system to address design issues, focusing on the optimal choice of the threshold.
Proposition 5.
Suppose Assumption 1 holds. Then, $V(R)$ is
  1. (i)
    decreasing in $R$ if $\sigma<1$;
  2. (ii)
    increasing in $R$ if $\sigma>1$;
  3. (iii)
    constant in $R$ if $\sigma=1$.
Intuitively, increasing $R$ makes the system more selective: fewer buy recommendations are issued, but those that are issued are more informative—i.e., they imply a higher expected payoff ($\Delta_{O}^{B}$). Conversely, lowering $R$ makes the system less selective, increasing the likelihood of buy recommendations at the cost of informativeness. When good products are more likely than bad ones ($\sigma>1$), it is optimal to be selective with buy recommendations, using a higher threshold to ensure that buy recommendations are more trustworthy. When bad products are more likely ($\sigma<1$), the opposite is true: being selective with don’t-buy recommendations is more valuable, and the platform should lower $R$. These forces exactly balance out when $\sigma=1$, so that the value of the system becomes independent of the threshold.
Put differently, the designer chooses whether the recommendation system should focus on identifying good or bad products. The threshold is set to reveal the more likely of these with greater certainty, aligning the selectiveness of the system with the underlying distribution of product quality.
4.2 Demand-Maximizing Recommendation Systems
We now turn to the analysis of demand-maximizing recommendation systems. Demand maximization requires a lenient recommendation threshold:
Proposition 6.
Suppose Assumption 1 holds. Then, demand is decreasing in $R$.
By Proposition 4 all types accept recommendations at all thresholds $R$ under Assumption 1. Thus, we can ignore the components of demand in equation (8) that come from receivers who do not accept don’t-buy recommendations. As a result, demand simplifies to $D(R)=\pi_{B}(R)$, and the result follows because the probability of buy recommendations is decreasing in $R$ [^16] See Lemma A.2 in the Appendix.
This result contrasts with Proposition 5 on value maximization. When $\sigma<1$, both value and demand are decreasing in $R$, so the objectives align. But when $\sigma>1$, they are diametrically opposed: while value increases with $R$, demand continues to fall. The underlying tradeoff is that increasing $R$ makes the system more selective—it raises the bar for a buy recommendation. From a value-maximization perspective, this can be desirable: fewer but more informative recommendations help receivers avoid bad matches, increasing their expected payoff. In this sense, selectivity enhances informativeness and thereby the value of the system.
By contrast, from a demand perspective, greater selectivity is never beneficial. The goal is simply to maximize the number of purchases, which is achieved by issuing buy recommendations as frequently as possible. A lenient system with a low threshold maximizes demand. Crucially, the symmetry assumption ensures that this leniency does not backfire: all types interpret recommendations in the same way and accept them. Thus, a low threshold increases demand without causing any mismatch between recommendation content and receiver preferences.
4.3 Preference Heterogeneity and Controversial Products
As discussed at the end of Section 3.1, preference heterogeneity only matters when controversial products exist. This naturally raises the question: does increasing heterogeneity—or increasing the share of controversial products—make recommendation systems less valuable? One might expect so, but the answer turns out to be more nuanced. To streamline the exposition, we focus only on value rather than demand in the two following subsections.
4.3.1 Preference Heterogeneity
To study the effect of changes in the preference heterogeneity, we consider mean-preserving spreads of the type distribution $F$, which preserve symmetry but increase the share of extreme types. It is useful to define the quantity $\beta:=F(1/2-R)$, which captures the fraction of the population willing to recommend a controversial product—either $(1,0)$ or $(0,1)$—at threshold $R$. For a fixed type distribution $F$, $\beta$ and $R$ are inversely related. Hence, we can interpret changes in $\beta$ as resulting either from changes in $R$ for a fixed distribution or from changes in the distribution itself at a fixed threshold. In what follows, we treat $\beta$ as a measure of polarization: for $R>1/2$, an increase in $\beta$ reflects a shift toward more extreme preferences, whereas for $R<1/2$, polarization corresponds to a decrease in $\beta$.
Proposition A.1 in Appendix A.2 shows formally how the value of the recommendation system depends on the threshold $R$ and the odds ratio $\sigma$. In the following, we sketch the main insights. Recall that $\beta$ represents the share of senders who are willing to recommend a controversial product. When $R>1/2$, a buy recommendation for a controversial product requires relatively strong preferences. [^17] An analogous argument applies to $R<1/2$. A higher $\beta$ then corresponds to an increase in polarization, making buy recommendations more frequent but less informative. Whether it is desirable to resolve this trade-off by increasing $\beta$ depends on $\sigma$: If good products are more likely than bad ones ($\sigma>1$), the system gains value from having buy recommendations that are highly informative, even if they are infrequent. In that case, polarization is harmful because it makes buy recommendations more common but more noisy. Conversely, when bad products are more likely ($\sigma<1$), the value of the system comes from being able to rule out bad products through lenient buy recommendations, and polarization helps by increasing their frequency.
4.3.2 Prevalence of Controversial Products
Without controversial products, preference heterogeneity would not matter: all receivers would accept recommendations, as these would reveal whether the product is good or bad. Beyond this extreme case, a natural conjecture is that recommendations become less valuable as controversial products become more common. We find that this conjecture does not hold without qualification.
When good and bad products are similarly likely, a higher prevalence of controversial products indeed reduces the value of the recommendation system. But when one type of product is much more likely than the other, the value of the system may actually be maximized at an interior share of controversial products. A formal statement of this result is given in Corollary 4 in Appendix A.2.
The reason lies in the structure of the value expression, which is the product of the buying probability and the objective informativeness of a buy recommendation (see Proposition 4). As controversial products become more prevalent, good products become less likely, and fewer buy recommendations are triggered by objectively good products. At the same time, more buy recommendations stem from controversial products. Thus, both the frequency and the informativeness of buy recommendations change. Whether the overall value increases or decreases depends on which of these effects dominates and on the odds-ratio $\sigma$. When good products are relatively rare ($\sigma<1$), identifying and avoiding bad products becomes especially valuable, which can raise the value of the system despite a lower share of objectively good recommendations. Conversely, when bad products are rare ($\sigma>1$), the value can increase as the recommendation system becomes more effective at identifying good ones, even if some recommendations are noisy.
5 Asymmetric Populations
While providing a useful benchmark, the assumption of a symmetric population does not allow us to study the role of the subjective component for the optimal design of the recommendation system. To make headway in that direction, we study asymmetric populations in this section. For the sake of tractability, we make the following assumption.
Assumption 2.
(i) $F(i)=(i+1/2)^{a}$ for $a>0$; (ii) $q_{1}=q_{2}=Q$.
We thus capture asymmetry by means of a simple but flexible parameterization. By assuming that the controversial products are equally likely a priori, we ensure that the subjective effect of a buy recommendation depends entirely on the asymmetry parameter $a$. When the type distribution is symmetric ($a=1$), there is no systematic bias in how controversial products are evaluated. As a result, buy recommendations do not shift beliefs about which controversial version the product might be—there is no subjective effect. In contrast, when the population is asymmetric (i.e., $a\neq 1$), one controversial version is systematically preferred over the other. In that case, buy recommendations tend to come from people who favor the majority-preferred version, and this introduces a systematic bias: The recommendation carries subjective information, shifting beliefs towards the more popular version.
5.1 Value-Maximizing Recommendation Systems
We can now analyze the extent to which the result in Proposition 5 that, generically, extreme thresholds are desirable is an artifact of the symmetry assumption. This is challenging because, contrary to the symmetric case, the set of types that accept recommendations depends on the threshold $R$. The following two results provide some insights into the optimal recommendation system in the case of an asymmetrically distributed population.
Proposition 7.
Suppose Assumption 2 holds. An optimal threshold $R^{\ast}\in(0,1)$ exists if $Q>1-\max\{\frac{a}{a+1},\frac{1}{a+1}\}$ or if (i) or (ii) holds:
[…]
Figure 2: Illustration of Propositions 7 and 8
[…]
Proposition 7, illustrated in Figure 2, provides sufficient conditions for interior solutions. It shows that an interior optimum exists when (i) controversial products are likely ($Q$ is large) and the population is very asymmetric ($a$ is far away from $1$) or otherwise when (ii) the uncontroversial product versions are similarly likely ($\sigma$ is close to $1$). [^18] The figure might suggest that the parameter region where $R^{\ast}\rightarrow 1$ is in some sense larger than the area where $R^{\ast}\rightarrow 0$. This is an artefact of the otherwise quite convenient parameterization of the preference distribution but does not have any significant economic meaning. The result starkly contrasts the symmetric case, where an interior solution is only possible in the knife-edge case $\sigma=1$. In that case, recommendations do not reveal any information about the relative likelihood of the controversial products beyond the prior, so the subjective payoff effect cancels out and the optimal threshold depends only on the odds ratio $\sigma$ of the uncontroversial products. In contrast, the average subjective payoff effect does not cancel out in the asymmetric case. Thus, the optimal threshold no longer only depends on $\sigma$ but also on the asymmetry parameter $a$ and the prevalence of controversial products $Q$. We explore under which circumstances the conditions given in Proposition 7 are not only sufficient but also necessary.
Proposition 8.
Suppose Assumption 2 holds. If $Q$ is small enough, then
  1. (i)
    $V(R)$ is decreasing in $R$ if $\sigma\leq\min\left\{\frac{1-Q(a+1)}{a-Q(a+1)},\frac{a-Q(a+1)}{1-Q(a+1)}\right\}$;
  2. (ii)
    $V(R)$ is increasing in $R$ if $\sigma\geq\max\left\{\frac{1-Q(a+1)}{a-Q(a+1)},\frac{a-Q(a+1)}{1-Q(a+1)}\right\}$;
  3. (iii)
    $V(R)$ is maximized for some $R^{\ast}\in(0,1)$ otherwise.
The proof proceeds in two steps. First, all types accept recommendations when $Q$ is sufficiently small, since then the subjective content in the recommendation is dominated by the objective content. This simplifies $V(R)$ by eliminating the case distinctions. We can use this information to prove the quasiconcavity of the function, ensuring that the sufficient conditions in Proposition 7 are necessary, too.
The result clarifies when interior thresholds emerge despite the recommendation system being inherently coarse. In the symmetric case, where subjective preferences cancel out on average, value maximization requires choosing an extreme threshold—either very lenient or very selective—depending on which uncontroversial product is more likely. The goal is to make recommendations essentially objective. However, in an asymmetric population, recommendations systematically shift beliefs about controversial products in favor of the majority-preferred version. This subjective component can be informative, especially when controversial products are common and the population is largely aligned in which one they prefer. In that case, even a buy recommendation for a controversial product is not very controversial—it likely signals the version that most people want. As a result, it can be optimal to back off from extreme thresholds and instead choose an intermediate threshold that allows drawing inferences about the controversial products, too. When preferences are less aligned or controversial products are rare, this logic breaks down, and more extreme thresholds become optimal again.
We next discuss how the interior threshold $R^{\ast}$—when it exists—responds to changes in the key parameters of the model. [^19] Due to the intractability of the problem, all of these statements are based on numerical simulations, see Appendix B. First, $R^{\ast}$ is increasing in $\sigma$: as good products become more likely relative to bad ones, the recommendation system should be more selective, and a higher threshold becomes optimal. Second, the further the preference distribution departs from symmetry (i.e., the further $a$ is from 1), the more $R^{\ast}$ moves to the interior. Intuitively, as $a$ moves further away from 1, the population becomes more homogeneous in its ranking of controversial products, so that including subjective content into the recommendation is informative and not too controversial. Third, the effect of the prevalence of controversial products $Q$ is more subtle. As $Q$ decreases, the setting effectively approaches that of a symmetric population, since asymmetry becomes irrelevant in the absence of controversial products. Accordingly, $R^{\ast}$ moves closer to the extreme threshold that would be optimal in the symmetric case: it approaches 1 if $\sigma>1$ and 0 if $\sigma<1$.
5.2 Demand-Maximizing Recommendation Systems
Like value maximization, demand maximization becomes significantly more complex when the population is asymmetrically distributed. Nonetheless, we establish a sufficient condition under which a low recommendation threshold is optimal, mirroring the result in the symmetric case.
Proposition 9.
Suppose $q_{H}\geq 1/2$. Then, demand is maximized for $R\rightarrow 0$.
While this result holds without Assumption 1 or Assumption 2, it restricts the analysis to the case that high quality goods are quite common. This complication arises because, contrary to the symmetric case, raising $R$ does not simply reduce the probability of buy recommendations—it also changes who follows them: As $R$ increases, controversial products are less likely to be recommended, and the composition of receivers who accept recommendations shifts toward those whose preferences are more aligned with the sender.
To understand the intuition for Proposition 9, recall that demand can come from two sources: types who accept buy recommendations and types who reject don’t-buy recommendations and choose to buy the product anyway. This suggests two potential strategies for maximizing demand. One option is to set $R$ low, ensuring that many buy recommendations are issued and that all types follow them (see Proposition 2). Another option would be to set an intermediate $R$ where not all types accept recommendations, thus relying on many people rejecting don’t-buy recommendations so that demand remains high even when buy recommendations are rarer. However, when $q_{H}\geq 1/2$, buy recommendations are always more frequent than don’t-buy recommendations. Consequently, the second approach is always dominated by the first approach. Thus, the optimal strategy is to maximize the frequency of buy recommendations by setting $R\to 0$.
This result stands in stark contrast with value maximization under asymmetry. According to Proposition 8, choosing as low a threshold as possible is optimal if and only if $\sigma\leq\min\left\{\frac{1-Q(a+1)}{a-Q(a+1)},\frac{a-Q(a+1)}{1-Q(a+1)}\right\}$ holds. However, given $q_{H}\geq 1/2$, it can be shown that this condition is never satisfied. This means that the objectives of demand and value maximization are never aligned in this setting: while demand is always maximized at $R\to 0$, value maximization typically favors a higher threshold.
Intuitively, note that good products are relatively frequent when $q_{H}\geq 1/2$, meaning that buy recommendations are already quite likely. In this case, the emphasis in value maximization shifts from frequency to informativeness: Since buy recommendations are common, their informativeness matters more than their frequency. Increasing $R$ filters out more controversial products, making buy recommendations more selective and thus more valuable. In contrast, demand maximization focuses exclusively on making buy recommendations as frequent as possible.
As the requirement that $q_{H}\geq 1/2$ is quite strong, we carry out a complementary numerical analysis (see Appendix B) for the case where it does not hold. For this analysis, we need to invoke Assumption 2. The results suggest that demand is monotonically decreasing in $R$, just as in the symmetric case. Therefore, as in the case of a symmetric population, demand maximization will still be aligned with value maximization for some parameters but not for others.
6 Robustness and Further Design Considerations
We now relax the assumptions that the preference distribution of senders and receivers is identical, that there are only two levels of recommendations (“buy” and “don’t-buy”) and that each receiver only has access to a single recommendation.
6.1 Distinct Sender and Receiver Distributions
Allowing for distinct sender and receiver distributions is interesting for two reasons. First, it captures situations such as when early adopters differ systematically from the broader population (similar to Acemoglu et al. (2022)). Second, and more importantly, it may be possible to shape the sender population or influence how receivers perceive it.
We first show that our earlier results extend naturally to the case of distinct sender and receiver distributions. Let the sender be drawn from $F$ and the receiver from $G$. If $F$ satisfies Assumption 1 or 2, the results on value- and demand-maximizing thresholds from Sections 4 and 5 remain valid. For example, Proposition A.2 in the appendix generalizes Proposition 5, showing that all receiver types accept recommendations under symmetry of $F$, and that $V(R)$ remains monotonic in $R$ with the slope depending on $\sigma$. Similarly, the potential conflict between value and demand objectives persists.
Beyond such robustness results, differing sender and receiver distributions give rise to new design considerations. In many cases, platforms may have partial control over the sender population or the information receivers have about it. For instance, Amazon’s phrasing “customers who bought this item also bought…” implicitly informs the receiver about who the senders are. We outline different ways to model this.
First, suppose the designer can manipulate the informativeness of a symmetric sender distribution $F$ through mean-preserving spreads or contractions. Though it may seem that making the sender distribution more precise would raise the value of the system, this is not always true. As in Section 4.3.1, changes in precision translate into changes in $\beta=F(1/2-R)$. With mild adaptations, the proof of Proposition A.2 can be used to show how the effect on value depends on the overall environment.
Second, suppose that the sender and receiver distributions are fixed, but that the designer can provide the receivers with a signal about the sender’s type. To make the point as starkly as possible, suppose that the signal fully reveals the type. If the sender is revealed to be $i=0$, all types accept recommendations, as the latter contain only objective information. In contrast to Proposition 2, where this behavior requires extreme thresholds, this would arise for all thresholds $R\in(0,1)$. Intuitively, as $i=0$ is indifferent between controversial products, her recommendation does not reveal subjective information. In contrast, if the signal reveals a sender of type $i\neq 0$, the receiver will be able to learn about controversial products, too. [^20] For instance, if $i=1$ and $R\in(0,1)$, a buy recommendation reveals that the product is either of version $(1,1)$ or $(1,0)$ Thus, having a sender who is partial to some controversial product is not necessarily a bad thing.
Third, suppose sender and receiver types are correlated so that recommendations are “tailored” to the receiver. In this case, the sender distribution becomes type-dependent, $F(i|\bar{i})$, and induces posterior beliefs that differ across receivers. While a suitably modified version of Proposition 1 still ensures that either all or no types follow recommendations, the set of types who do so now depends on the joint distribution. This leads to complications in further analyses similar to those discussed in Section 5 for the case of asymmetric populations.
6.2 Multiple Recommendation Thresholds
We now consider a recommendation system with finer granularity. Specifically, we relax the benchmark assumption of binary recommendations by introducing a third, intermediate option. Such richer recommendation structures appear in real-world platforms and are also shown to be optimal in models of strategic communication under social learning (e.g., Smirnov and Starkov, 2024; Bénabou and Vellodi, 2024).
Formally, we introduce two threshold levels $R_{1},R_{2}\in(0,1)$ with $R_{2}\geq R_{1}$, and allow for three recommendation types $r\in\{D,N,B\}$: “don’t-buy,” “neutral,” and “buy.” The sender issues a buy recommendation if her payoff is at least $R_{2}$, a don’t-buy if it is below $R_{1}$, and a neutral recommendation otherwise. We maintain Assumption 1 throughout.
As in the benchmark model, the highest and lowest recommendations convey objective information: $r=B$ rules out bad products, and $r=D$ rules out good ones. The neutral recommendation, by contrast, carries no objective content; it implies that the product is controversial, but does not distinguish between versions $(1,0)$ and $(0,1)$. This leads to the following result:
Proposition 10.
Suppose there are recommendation thresholds $R_{2}\geq R_{1}$ and Assumption 1 holds. Then $V(\mathcal{R})$ is (i) decreasing in $R_{1}$ if $\sigma<1$, and (ii) increasing in $R_{2}$ if $\sigma>1$.
This result mirrors the logic of Proposition 5: the designer maximizes the value of the recommendation system by ensuring that one of the thresholds produces recommendations that are essentially objective. The optimal threshold depends on which uncontroversial product version is more likely. If, for instance, bad products are more prevalent ($\sigma<1$), the low threshold $R_{1}$ should be set close to zero to ensure that don’t-buy recommendations are informative. The optimal value of the remaining threshold depends on the model parameters, unless $q_{1}=q_{2}$, in which case any choice yields the same outcome. In some cases, both thresholds may be set at extreme values, rendering both $r=B$ and $r=D$ essentially objective. [^21] This arises, for instance, under $F(i)=1/2+i$ with $q_{1}=0.7$, $q_{2}=0.1$, and $q_{H}=0.03$.
For a demand-maximizing platform, however, richer recommendation systems offer no benefit under Assumption 1. Since receivers of the neutral recommendation are only partially informed, they may or may not buy the product, depending on the environment. As shown in the proof of Proposition 10, this ambiguity leads to weakly lower demand than under a binary system with a very low threshold $R\to 0$, which maximizes the number of buy recommendations. Hence, the intermediate recommendation dampens demand without offsetting gains in platform revenue.
6.3 Multiple Recommendations
We now show that the value of a recommendation system does not always increase if receivers obtain more than one recommendation: The additional recommendations do not add value if the ex ante probabilities of the good and bad products are similar while those of the two controversial product versions are not.
Formally, suppose the receiver observes $b$ buy and $d$ don’t-buy recommendations, each independently drawn from a population of senders as in the benchmark model. [^22] We derive the posteriors for general $(b,d)$ in Appendix A.5. A single buy recommendation is sufficient to rule out a bad product, and a single don’t-buy recommendation rules out a good product. Hence, when the receiver obtains a mix of both, he infers that the product is not uncontroversial—it must be one of the two controversial versions. [^23] See also Sun (2011), who relate recommendation variance to product “nicheness” in a similar spirit. To highlight the limits of informational gains from additional recommendations, we focus on the case of infinite learning, where the receiver observes an unbounded number of independent recommendations.
Proposition 11.
Consider the model with multiple recommendations. Suppose Assumption 1 holds and let $\lambda:=q_{1}/q_{2}$. Then, infinite learning does not increase the value of the recommendation system relative to the optimally designed single-recommendation benchmark if and only if one of the following holds:
[…]
The intuition is as follows. With infinitely many recommendations, the receiver can perfectly infer whether the product is good, bad, or controversial. As in the benchmark case, receivers buy good products and reject bad ones. Thus, any potential gain must come from improved decisions about controversial products. However, controversial products always generate mixed recommendations. Whether the receiver ultimately chooses to buy depends on his posterior belief and thus on the relative likelihood of the outside option being good or bad. If good products are relatively rare (that is, $\sigma$ is low), the outside option is unattractive, and the receiver may prefer a controversial product even if it is only an imperfect match. If good products are common (that is, $\sigma$ is high), the receiver is more willing to choose the outside option. In either extreme, the decision under infinite learning coincides with what could already be achieved by a single recommendation from an optimally calibrated recommendation system. Only when the uncontroversial versions are similarly likely and controversial products are sufficiently ambiguous does additional information add value.
Finally, note that providing more recommendations also does not help a demand-maximizing platform under Assumption 1. If the platform’s goal is to induce as many purchases as possible, it suffices to choose a single threshold $R\rightarrow 0$, which leads all senders to issue buy recommendations. Since mixed recommendations can only reduce the likelihood of purchasing controversial products, adding them weakly reduces demand.
7 Conclusion
This paper has examined under which circumstances coarse, non-strategic recommendations can meaningfully guide decision-making in environments with heterogeneous preferences and subjective product quality. Our analysis highlights that even minimal feedback, such as a binary “likes”, can transmit valuable information.
At the heart of our approach is the distinction between the objective and subjective content of a recommendation. While a buy or don’t-buy signal always reveals something about product quality, its usefulness depends on how much of that signal reflects shared preferences versus individual taste. We show that when recommendations are largely objective, all receivers are willing to follow them, either by construction (e.g., through extreme thresholds) or by aggregation (e.g., in symmetric populations).
Systematic preference heterogeneity complicates this picture. In asymmetric populations, where controversial products are common and preferences are aligned in one direction, a recommendation typically contains both objective and subjective content. The same signal—say, a buy recommendation—may reflect a product’s broad appeal, or it may simply reveal that it aligns with the preferences of a particular sender. In such settings, subjectivity introduces friction: Receivers must infer how much of the recommendation reflects shared values rather than idiosyncratic taste, and in some cases, rejecting a recommendation may be the better choice.
However, subjectivity need not be treated solely as a source of noise. When preference alignment is strong and controversial products are sufficiently prevalent, many controversial products are in fact not very controversial in the eyes of the majority. Then, an optimally designed recommendation system enables meaningful inference about such products by preserving some subjective content. In contrast, when this condition fails, it becomes optimal to adopt an extreme threshold that filters out all subjective noise. This stands in contrast to the symmetric population case, where the subjective content of the recommendation system should be suppressed by focusing entirely on extreme thresholds.
Throughout, we contrast the objectives of maximizing user value with maximizing demand. These goals are sometimes aligned, but they often diverge. A platform seeking to drive demand may prefer lenient recommendation thresholds that generate more buy signals, even if this reduces informativeness and harms users.
Finally, we have maintained a number of classical assumptions such as risk neutrality, linear probability weights, and Bayesian updating. Preliminary results suggest that a promising avenue for future research is to relax these assumptions so as to study their impact on the receiver’s tendency to accept a recommendation or not and, thus, on the value of the recommendation system.
Appendix A Formal Details and Proofs
A.1 Proofs and Technical Details for the Benchmark Model
Posterior Distributions
The receiver’s posterior $\mathbf{p}^{B}(R):=(p_{H}^{B}(R),p_{1}^{B}(R),p_{2}^{B}(R),p_{L}^{B}(R))$ after a buy recommendation is
$p_{H}^{B}(R):=\frac{q_{H}}{q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R)},\quad\quad$ $p_{2}^{B}(R):=\frac{q_{2}\phi_{2}(R)}{q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R)},$
$p_{1}^{B}(R):=\frac{q_{1}\phi_{1}(R)}{q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R)},\quad\quad$ $p_{L}^{B}(R):=0.$
A don’t-buy recommendation gives rise to $\mathbf{p}^{D}(R):=(p_{H}^{D}(R),p_{1}^{D}(R),p_{2}^{D}(R),p_{L}^{D}(R))$, where
$p_{H}^{D}(R):=0,\,$ $p_{2}^{D}(R):=\frac{q_{2}(1-\phi_{2}(R))}{q_{1}(1-\phi_{1}(R))+q_{2}(1-\phi_{2}(R))+q_{L}},$
$p_{1}^{D}(R):=\frac{q_{1}(1-\phi_{1}(R))}{q_{1}(1-\phi_{1}(R))+q_{2}(1-\phi_{2}(R))+q_{L}},\,$ $p_{L}^{D}(R):=\frac{q_{L}}{q_{1}(1-\phi_{1}(R))+q_{2}(1-\phi_{2}(R)))+q_{L}}.$
Proof of Proposition 1
Part (i): Let $\Delta^{B}(i)\equiv U^{B}(i)-U^{0}(i)$ be the gain from following a buy–recommendation and $\Delta^{D}(i)\equiv U^{0}(i)-U^{D}(i)$ the gain from following a don’t-buy–recommendation. A receiver follows a recommendation iff the corresponding gain is non–negative.
For every state $s\in\{H,1,2,L\}$, Bayes’ rule implies
$\pi^{B}(p^{B}_{s}-q_{s})+\pi^{D}(p^{D}_{s}-q_{s})=0.$
Multiplying by the state–contingent payoffs $\{1,1/2+i,1/2-i,0\}$ and summing over the states we obtain
$\pi^{B}\Delta^{B}(i))=\pi^{D}\Delta^{D}(i).$
Because $\pi^{B},\pi^{D}>0$, the differences $\Delta^{D}(i)>0$ and $\Delta^{D}(i)$ have the same sign, so that the receiver either follows both recommendations or rejects both, proving part (i). Part (ii): This follows directly from plugging in $\Delta^{r}_{O}$ and $\Delta^{r}_{S}$ into equation (1) for the buy recommendation the analogous equation for a don’t-buy recommendation.
Proof of Proposition 2
Recall from equation (1) the optimal acceptance condition
$(p_{H}^{B}-q_{H})+\frac{p_{1}^{B}-q_{1}}{2}+\frac{p_{2}^{B}-q_{2}}{2}\geq i[(p_{2}^{B}-q_{2})-(p_{1}^{B}-q_{1})].$
First, observe that
$\lim_{R\rightarrow 1}p_{1}^{B}=\lim_{R\rightarrow 1}p_{2}^{B}=0$
and thus $\lim_{R\rightarrow 1}p_{H}^{B}=1$. Hence, in the case $R\rightarrow 1$ the optimality condition simplifies to
$1-q_{H}-\frac{q_{1}+q_{2}}{2}\geq i(q_{1}-q_{2}).$ (9)
Equation (9) always holds for $i\geq 0$. Further, for $q_{1}>q_{2}$ the condition is most stringent for $i=1/2$. In this case, it is equivalent to $1\geq q_{1}+q_{H}$, which always holds. The case $i<0$ is analogous. In the case $R\rightarrow 0$ we get
$\lim_{R\rightarrow 0}p_{s}(B)=\frac{q_{s}}{1-q_{L}}$
for $s\in\{H,1,2\}$ and thus the optimality condition simplifies to
$\frac{q_{H}q_{L}}{1-q_{L}}+\frac{q_{1}q_{L}}{2(1-q_{L})}+\frac{q_{2}q_{L}}{2(1-q_{L})}\geq i[\frac{q_{2}q_{L}}{1-q_{L}}-\frac{q_{1}q_{L}}{1-q_{L}}]$
$\Leftrightarrow 2q_{H}+q_{1}+q_{2}\geq 2i[q_{2}-q_{1}].$
For $i\geq 0$, the last condition always holds if $q_{1}\geq q_{2}$. Further, for $q_{2}>q_{1}$ the condition is most stringent for $i=1/2$. In this case, it is equivalent to $q_{H}+q_{1}\geq 0$, which is always satisfied. The case $i<0$ is analogous.
Proof of Proposition 3
(i) If $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, inserting (5) and (6) yields
[…]
$=\pi^{B}\int_{-1/2}^{1/2}\bigl{(}p_{H}^{B}+(1/2+i)p_{1}^{B}+(1/2-i)p_{2}^{B}\bigr{)}-(q_{H}+(1/2+i)q_{1}+(1/2-i)q_{2})dF(i)$
[…]
$=\pi^{B}[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]]$
(ii) If $\Delta_{S}^{B}<-2\Delta_{O}^{B}$, inserting (5), (6) and (7) yields
[…]
$=\int_{-1/2}^{\tilde{i}}\pi^{B}\left(q_{H}+q_{1}(\frac{1}{2}+i)+q_{2}(\frac{1}{2}-i)\right)+(1-\pi^{B})\left(p_{H}^{D}+p_{1}^{D}(\frac{1}{2}+i)+p_{2}^{D}(\frac{1}{2}-i)\right)dF(i)$
$+\int_{\tilde{i}}^{1/2}\pi^{B}\left(p_{H}^{B}+p_{1}^{B}(\frac{1}{2}+i)+p_{2}^{B}(\frac{1}{2}-i)\right)+(1-\pi^{B})\left(q_{H}+q_{1}(\frac{1}{2}+i)+q_{2}(\frac{1}{2}-i)\right)dF(i)$
$-\int_{-1/2}^{\tilde{i}}q_{H}+q_{1}(\frac{1}{2}+i)+q_{2}(\frac{1}{2}-i)dF(i)-\int_{\tilde{i}}^{1/2}q_{H}+q_{1}(\frac{1}{2}+i)+q_{2}(\frac{1}{2}-i)dF(i)$
$=(1-\pi^{B})\int_{-1/2}^{\tilde{i}}\underbrace{(p_{H}^{D}-q_{H})+\frac{p_{1}^{D}-q_{1}}{2}+\frac{p_{2}^{D}-q_{2}}{2}}_{=\Delta_{O}^{D}}-i(\underbrace{[(p_{2}^{D}-q_{2})-(p_{1}^{D}-q_{1})]}_{=\Delta_{S}^{D}})dF(i)$
$+\pi^{B}\int_{\tilde{i}}^{1/2}\underbrace{(p_{H}^{B}-q_{H})+\frac{p_{1}^{B}-q_{1}}{2}+\frac{p_{2}^{B}-q_{2}}{2}}_{=\Delta_{O}^{B}}-i(\underbrace{[(p_{2}^{B}-q_{2})-(p_{1}^{B}-q_{1})]}_{=\Delta_{S}^{B}})dF(i)$
$=(1-\pi^{B})F(\tilde{i})\left[\Delta_{O}^{D}-\Delta_{S}^{D}\mathbb{E}[i\mid i\leq\tilde{i}]\right]+\pi^{B}(1-F(\tilde{i}))\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i\mid i\geq\tilde{i}]\right].$
[…]
Proof of Proposition 4
According to Corollary 1, all types accept the recommendation if
$\min\{\Delta_{O}^{B}+\frac{\Delta_{S}^{B}}{2},\Delta_{O}^{B}-\frac{\Delta_{S}^{B}}{2}\}\geq 0.$ (10)
To see that this condition is fulfilled, first observe that
$\Delta_{O}^{B}+\frac{\Delta_{S}^{B}}{2}$ $=p_{H}^{B}-q_{H}+p_{2}^{B}-q_{2}$
$=\frac{q_{H}}{q_{H}+\beta(q_{1}+q_{2})}-q_{H}+\frac{\beta q_{2}}{q_{H}+\beta(q_{1}+q_{2})}-q_{2}$
$=\frac{q_{H}+\beta q_{2}-(q_{H}+q_{2})(q_{H}+\beta(q_{1}+q_{2}))}{q_{H}+\beta(q_{1}+q_{2})}$
$=\frac{\beta(q_{2}-(q_{H}+q_{2})(q_{1}+q_{2}))+q_{H}(1-q_{H}-q_{2})}{q_{H}+\beta(q_{1}+q_{2})}$
This term is positive if $q_{2}-(q_{H}+q_{2})(q_{1}+q_{2})\geq 0$. If $q_{2}-(q_{H}+q_{2})(q_{1}+q_{2})<0$, then
$\frac{\beta(q_{2}-(q_{H}+q_{2})(q_{1}+q_{2}))+q_{H}(1-q_{H}-q_{2})}{q_{H}+\beta(q_{1}+q_{2})}$
$\geq\frac{(q_{2}-(q_{H}+q_{2})(q_{1}+q_{2}))+q_{H}(1-q_{H}-q_{2})}{q_{H}+\beta(q_{1}+q_{2})}=\frac{(q_{H}+q_{2})(1-q_{1}-q_{2}-q_{H})}{q_{H}+\beta(q_{1}+q_{2})}\geq 0\text{.}$
Thus, $\Delta_{O}^{B}+\frac{\Delta_{S}^{B}}{2}\geq 0$ even in this case. Proceeding analogously, one obtains $\Delta_{O}^{B}-\frac{\Delta_{S}^{B}}{2}=p_{H}^{B}-q_{H}+p_{1}^{B}-q_{1}\geq 0.$ Corollary 1 thus implies that all types accept the recommendation. Assumption 1 implies $\mathbb{E}[i]=0$, so that the result follows from Proposition 3.
Proof of Proposition 5
We first restate the probabilities of buy recommendations, their objective content and the value of the recommendation system in terms of $\sigma$, $\beta$ and $Q$:
Lemma A.1.
Suppose Assumption 1 holds. Then,
$\pi^{B}$ $=\frac{(1-2Q)\sigma}{1+\sigma}+2Q\beta;$
$\Delta_{O}^{B}$ $=\frac{\beta Q(\sigma+1)-2Q\sigma+\sigma}{2\beta Q(\sigma+1)-2Q\sigma+\sigma}-\frac{(1-2Q)\sigma}{1+\sigma}-Q;$
$V(\beta)$ $=\pi^{B}\Delta_{O}^{B}=(1-2Q)\frac{\sigma+Q\left(\beta-\sigma+\sigma^{2}(1-\beta)\right)}{(\sigma+1)^{2}}.$
Proof.
The result on $\pi^{B}$ follows from inserting $q_{H}=(1-2Q)\sigma/(1+\sigma)$, $\phi_{1}(R)=\phi_{2}(R)=\beta$ and $q_{1}+q_{2}=2Q$ into $\pi^{B}=q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R)$. Next, equation (2) gives
$\Delta_{O}^{B}=\frac{\beta Q(\sigma+1)-2Q\sigma+\sigma}{2\beta Q(\sigma+1)-2Q\sigma+\sigma}-\frac{(1-2Q)\sigma}{1+\sigma}-Q$ (11)
The result on $V(\beta)$ then follows immediately from Proposition 4.
The following result shows that countervailing effects of $\beta$ on the probability of a buy recommendation and the objective value $\Delta_{O}^{B}$ of accepting it result in a non-trivial effect on $V(\beta)=\pi^{B}\Delta_{O}^{B}$.
Lemma A.2.
Suppose Assumption 1 holds. Then,
  • (i)
    $\pi^{B}$ is strictly increasing and $\Delta_{O}^{B}$ strictly decreasing in $\beta$;
  • (ii)
    $\frac{\partial^{2}V}{\partial\beta\partial\sigma}<0$ and $V^{\prime}(\beta)=\begin{cases}>0&\text{if }\sigma<1;\\ =0&\text{if }\sigma=1;\\ <0&\text{if }\sigma>1.\end{cases}$
Proof.
(i) Note that $\pi^{B}=(1-2Q)\sigma/(1+\sigma)+2Q\beta$, so that $\pi_{B}$ is strictly increasing. Moreover,
$\frac{\partial{\Delta_{O}^{B}}}{\partial{\beta}}=-\frac{(Q(1-2Q)\sigma(1+\sigma))}{(\sigma-2Q\sigma+2\beta Q(1+\sigma))^{2}}<0.$
(ii) Lemma A.1 implies
$V^{\prime}(\beta)=\frac{Q(1-2Q)(1-\sigma)}{1+\sigma},$
and hence
$\frac{\partial{{}^{2}V}}{\partial{\beta\partial\sigma}}=-\frac{2Q(1-2Q)}{(1+\sigma)^{2}}<0,$
$V^{\prime}(\beta)$ is negative for $\sigma>1$, positive for $\sigma<1$ and zero for $\sigma=1$.
Then, using $\beta=F(1/2-R)$, Lemma A.2 directly implies Proposition 5.242424Result (iii) follows because $V(\beta)$ is independent of $\beta$ in this case.
Proof of Proposition 7
[…]
$V(R)=\begin{cases}\pi^{B}\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]\right]&\text{if }\tilde{i}>1/2\\ \pi^{B}F(\tilde{i})\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i\mid i\leq\tilde{i}]\right]&\text{if }\tilde{i}\leq 1/2\\ +(1-\pi^{B})(1-F(\tilde{i}))\left[\Delta_{O}^{D}-\Delta_{S}^{D}\mathbb{E}[i\mid i\geq\tilde{i}]\right]&\\ \end{cases}$
Proposition 2 implies $\tilde{i}>1/2$ if $R\rightarrow 0$ or $R\rightarrow 1$. Thus, it suffices to consider
$V(R)=$ $(q_{H}+Q(1-R^{a}+(1-R)^{a}))\left(\frac{2q_{H}+Q(1-R^{a}+(1-R)^{a})}{2(q_{H}+Q(1-R^{a}+(1-R)^{a}))}-q_{H}-Q\right)$
$+(q_{H}+Q(1-R^{a}+(1-R)^{a}))\frac{Q(1-R^{a}-(1-R)^{a})}{q_{H}+Q(1-R^{a}+(1-R)^{a})}\frac{a-1}{2(a+1)}$
$=\left(\frac{2q_{H}+Q(1-R^{a}+(1-R)^{a})}{2}-(q_{H}+Q(1-R^{a}+(1-R)^{a}))(q_{H}+Q)\right)$
$+Q(1-R^{a}-(1-R)^{a})\frac{a-1}{2(a+1)}$
for $R$ sufficiently close to 0 and 1. Using $q_{H}=\frac{(1-2Q)\sigma}{1+\sigma}$ and rearranging gives
$V(R)$ $=\frac{(1-2Q)\sigma}{1+\sigma}+\frac{aQ(1-R^{a})+Q(1-R)^{a}}{a+1}$
$-\left(\frac{(1-2Q)\sigma}{1+\sigma}+Q(1-R^{a})+Q(1-R)^{a}\right)\left(\frac{(1-2Q)\sigma}{1+\sigma}+Q\right)$
We then obtain the corresponding derivative
$V^{\prime}(R)=\frac{aQ\left((a+1)\left(R^{a-1}+(1-R)^{a-1}\right)(Q+\frac{(1-2Q)\sigma}{1+\sigma})-aR^{a-1}-(1-R)^{a-1}\right)}{a+1}$
We can then write
$V^{\prime}(R)>0$
$\Leftrightarrow(a+1)\left(R^{a-1}+(1-R)^{a-1}\right)(Q+\frac{(1-2Q)\sigma}{1+\sigma})-aR^{a-1}-(1-R)^{a-1}>0$
$\Leftrightarrow(a+1)\left(Q+\frac{(1-2Q)\sigma}{1+\sigma}\right)>\frac{aR^{a-1}+(1-R)^{a-1}}{R^{a-1}+(1-R)^{a-1}}=\frac{a(1-R)^{1-a}+R^{1-a}}{(1-R)^{1-a}+R^{1-a}},$
so that
$\lim_{R\rightarrow 0}V^{\prime}(R)>0\Leftrightarrow(a+1)\left(Q+\frac{(1-2Q)\sigma}{1+\sigma}\right)>a$
and analogously
$\lim_{R\rightarrow 1}V^{\prime}(R)<0\Leftrightarrow(a+1)\left(Q+\frac{(1-2Q)\sigma}{1+\sigma}\right)<1.$
Thus, a sufficient condition for an interior solution when $a<1$ is given by
$\frac{1}{a+1}>\left(Q+\frac{(1-2Q)\sigma}{1+\sigma}\right)>\frac{a}{a+1}.$ (12)
Rearranging the first inequality, we obtain
$\left(\frac{1}{a+1}-Q\right)(1+\sigma)>(1-2Q)\sigma$
$\Leftrightarrow\frac{1}{a+1}-Q>(1-2Q)\sigma-\left(\frac{1}{a+1}-Q\right)\sigma$
$\Leftrightarrow\frac{1}{a+1}-Q>\sigma\left(\frac{a}{1+a}-Q\right)$
For $a<1$, the condition $Q\geq 1-\max\{\frac{a}{a+1},\frac{1}{a+1}\}$ is equivalent to $Q\geq\frac{a}{1+a}$ so that the right-hand side is negative while the left-hand side is positive. Hence, this inequality is always satisfied. Proceeding analogously for the second inequality, we obtain
$\left(\frac{a}{a+1}-Q\right)(1+\sigma)<(1-2Q)\sigma,$
which is also implied by $Q\geq 1-\max\{\frac{a}{a+1},\frac{1}{a+1}\}$. Hence, the first statement of Proposition 7 for the case $a<1$ follows. For the second statement when $Q<1-\max\{\frac{a}{a+1},\frac{1}{a+1}\}$, we can rewrite the two inequalities in (12) as
$\frac{1-Q(a+1)}{a-Q(a+1)}>\sigma>\frac{a-Q(a+1)}{1-Q(a+1)}$
giving the second condition in the statement.
Proof of Proposition 8
Observe that
$\frac{\partial\tilde{i}}{\partial Q}=\frac{Q^{2}\left(R^{a}-(1-R)^{a}-1\right)+(q_{H}-1)q_{H}}{Q^{2}\left(R^{a}+(1-R)^{a}-1\right)}$ (13)
so that $\tilde{i}$ is monotonically increasing in $Q$ for $a>1$ and monotonically decreasing for $a<1$. Moreover, we have
$\lim_{Q\rightarrow 0}\Delta_{S}^{B}$ $=\lim_{Q\rightarrow 0}-\frac{Q(1-R^{a}-(1-R)^{a})}{q_{H}+Q(1-R^{a}+(1-R)^{a})}=0,$
and
$\lim_{Q\rightarrow 0}\Delta_{O}^{B}$ $=\lim_{Q\rightarrow 0}\frac{2q_{H}+Q(1-R^{a}+(1-R)^{a})}{2(q_{H}+Q(1-R^{a}+(1-R)^{a}))}-q_{H}-Q=1-q_{H}>0.$
Therefore, $\tilde{i}$ diverges as $Q\rightarrow 0$. Hence, by continuity and Corollary 1, there exists a cutoff $\bar{Q}$ below which all types accept recommendations. Thus, Proposition 3 implies
$V(R)$ $=\pi^{B}[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]]$
Rearranging this equation gives
$V(R)$ $=\frac{(1-2Q)\sigma}{1+\sigma}+\frac{aQ(1-R^{a})+Q(1-R)^{a}}{a+1}$
$-\left(\frac{(1-2Q)\sigma}{1+\sigma}+Q(1-R^{a})+Q(1-R)^{a}\right)\left(\frac{(1-2Q)\sigma}{1+\sigma}+Q\right)$
$=\frac{(1-2Q)\sigma}{1+\sigma}\left(1-\left(\frac{(1-2Q)\sigma}{1+\sigma}+Q\right)\right)$
$+\frac{Q}{a+1}(a(1-R^{a})+(1-R)^{a})$
$-\frac{Q}{a+1}\left((1-R^{a})+(1-R)^{a}\right)(a+1)\left(\frac{(1-2Q)\sigma}{1+\sigma}+Q\right)$
$=\frac{(1-2Q)\sigma(Q(\sigma-1)+1)}{(\sigma+1)^{2}}$
$+\frac{Q}{a+1}(1-R^{a})\left(a-(a+1)\left(\frac{(1-2Q)\sigma}{1+\sigma}+Q\right)\right)$
$+\frac{Q}{a+1}(1-R)^{a}\left(1-(a+1)\left(\frac{(1-2Q)\sigma}{1+\sigma}+Q\right)\right)$
$=c_{0}+\frac{Q}{a+1}\left[c_{1}(a)(1-R^{a})+c_{2}(a)(1-R)^{a}\right]$
with $c_{0}=\frac{(1-2Q)\sigma(Q(\sigma-1)+1)}{(\sigma+1)^{2}}$, $c_{1}(a)=a+(a+1)\frac{Q(\sigma-1)-\sigma}{\sigma+1}$ and $c_{2}(a)=1+(1+a)\frac{Q(\sigma-1)-\sigma}{\sigma+1}$. We thus obtain
$V^{\prime}(R)$ $=-\frac{Qa}{a+1}\left[R^{a-1}c_{1}(a)+(1-R)^{a-1}c_{2}(a)\right],$
$V^{\prime\prime}(R)$ $=-\frac{Qa(a-1)}{a+1}\left[R^{a-2}c_{1}(a)-(1-R)^{a-2}c_{2}(a)\right],$
For $c_{1}(a),c_{2}(a)\geq 0$ and $c_{1}(a),c_{2}(a)\leq 0$ the function $V(R)$ is monotonically decreasing and increasing, respectively. Further, for $a\geq 1$, $c_{1}(a)>0$ and $c_{2}(a)<0$ it is concave. Conversely, for $a<1$ and $c_{1}(a)<0$ and $c_{2}(a)>0$ it is concave. Further,
$c_{1}(a)-c_{2}(a)\begin{cases}<0&a<1\\ =0&a=1\\ >0&a>1,\end{cases}$
Therefore, for $a<1$ either (i) $c_{1}(a),c_{2}(a)\leq 0$, (ii) $c_{1}(a)\leq 0\leq c_{2}(a)$, or (iii) $0\leq c_{1}(a),c_{2}(a)$. Hence, in all three cases $V(R)$ is quasiconcave. Similarly, for $a>1$ either (i) $c_{1}(a),c_{2}(a)\leq 0$, (ii) $c_{2}(a)\leq 0\leq c_{1}(a)$, or (iii) $0\leq c_{1}(a),c_{2}(a)$. Hence, in all three cases $V(R)$ is quasiconcave. Finally, for $a=1$ quasiconcavity follows because $c_{1}(a)=c_{2}(a)$ and thus $V$ is monotone. Putting all of this together, it follows that $V$ is quasiconcave in $R$. Hence, the sufficient conditions in Proposition 7 for an interior solution are also necessary.
Proof of Proposition 9
We need to show that $D(R)\leq D(0)\forall R>0$. This is equivalent with
$\pi_{B}(R)\hat{F}(R)+(1-\pi_{B}(R))(1-\hat{F}(R))\leq q_{H}+q_{1}+q_{2}$ (14)
To show (14), first note that $q_{H}\geq 1/2$ implies that $\pi_{B}(R)=q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R)\geq 1/2$. Next, $\pi_{B}(R)\geq 1/2$ is equivalent with
$\pi_{B}(R)\hat{F}(R)+(1-\pi_{B}(R))(1-\hat{F}(R))\leq\pi_{B}(R)$ (15)
From $D(0)=q_{H}+q_{1}+q_{2}$, we further obtain $\pi_{B}(R)\leq D(0)$. Together, this gives (14).
A.2 Results from Section 4.3
Increasing Heterogeneity (Section 4.3.1)
The following result captures the effects of a mean-preserving spread on the value of the recommendation system.
Proposition A.1.
Suppose Assumption 1 holds.
  1. 1.
    If $\sigma<1$, then a mean-preserving spread of the type distribution increases the value of the recommendation system when $R>1/2$ and decreases it when $R<1/2$.
  2. 2.
    If $\sigma>1$, then a mean-preserving spread of the type distribution decreases the value of the recommendation system when $R>1/2$ and increases it when $R<1/2$.
Proof.
As argued in the text, a mean-preserving spread corresponds to an increase in $\beta$ for $R>1/2$, while for $R<1/2$ it corresponds to a decrease in $\beta$. From the proof of Lemma A.2, we can see that the value of the recommendation system is increasing in $\beta$ for $\sigma<1$ and decreasing in $\beta$ for $\sigma>1$. Thus, the statement follows.
To build intuition, suppose $\sigma<1$ and $R>1/2$. As shown in Figure 1, when $R>1/2$, only extreme types issue buy recommendations for controversial products. [^26] For $R<1/2$, only relatively extreme types issue don’t-buy recommendations for controversial products. This follows by reversing the colors in Figure 1 to correspond to $R=1/4$. A mean-preserving spread of $F$ (which corresponds to an increase in $\beta$ given $R$), thus leads to more extreme senders and thus a higher likelihood that buy recommendations stem from controversial products. This raises the probability of buy recommendations while lowering their objective value. The net effect, and whether the system’s value rises or falls, depends on the odds ratio $\sigma$ (see Proposition 5). When $\sigma<1$, good products are rarer than bad ones, making the system more valuable as $\beta$ increases. The benefit of revealing bad products by means of don’t-buy recommendations outweighs the reduced value of buy recommendations. Thus, for high $R$, a more polarized population benefits more from the system, as only bad products result in don’t-buy recommendations. Conversely, for low $R$, a less polarized population centered around the middle ensures that only bad products receive don’t-buy recommendations, thereby yielding a higher value.
Figure 3: Population polarization
[…]
Turning to the impact of preference heterogeneity on demand, the analysis is simpler. Recall that under Assumption 1, all receiver types accept recommendations. Consequently, it follows from Lemma A.2 that demand is increasing in $\beta$. Finally, the same distinction depending on whether $R>1/2$ or $R<1/2$ in terms of the impact of a mean-preserving spread applies as above.
The Prevalence of Controversial Products (Section 4.3.2)
We prove the following result:
Corollary A.1.
Suppose Assumption 1 holds. Then for every $\beta\in(0,1)$, there exist $\underline{\sigma}(\beta)\in(0,1)$ and $\overline{\sigma}(\beta)>1$ such that
  • (i)
    the value of the recommendation system is decreasing in $Q$ on $(\underline{\sigma}(\beta),\overline{\sigma}(\beta))$.
  • (ii)
    For $\sigma<\underline{\sigma}$ and $\sigma>\underline{\sigma}$, the value of the recommendation system is maximized for some $Q^{*}>0$.
Figure 4: Illustration of Corollary 4
(a) Panel A
$\sigma$$\beta$111133335555777799990.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81.01.01.01.0
(b) Panel B
$\sigma$$\beta$111133335555777799990.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81.01.01.01.0
(c) Panel C
$\sigma$$\beta$111133335555777799990.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81.01.01.01.0
Figure 4 illustrates the result. In the white area in Panel A, $V(\beta)$ is decreasing in $\beta$; in the shaded area, an interior maximizer exists. When good and bad products are similarly likely ($\sigma$ is close to $1$), a higher prevalence of controversial products indeed reduces $V(\beta)$. For very high or low values of $\sigma$, however, this need not be true. This surprising result reflects the multiplicative structure of $V(\beta)$. Under Assumption 1, $V(\beta)$ is the product of the buying probability $\pi^{B}$ and the objective value of the recommendation per buyer, $\Delta_{O}^{B}$. The effect of $Q$ on $\pi^{B}$ is negative only if $\sigma$ is sufficiently large (see Panel B). Intuitively, as $Q$ increases, good products are less likely, reducing the probability of buy recommendations. However, there are now more buy recommendations stemming from controversial products, as these have become more likely. $\beta$ and $\sigma$ determine which of these effects dominates. When $\beta$ is high, the greater likelihood that controversial products receive buy recommendations increases the total probability of buy recommendations, strengthening the positive effect. When $\sigma$ is high, the lower share of good products dominates, strengthening the negative effect. Similarly, one can show that the effect of $Q$ on $\Delta_{O}^{B}$ becomes negative only for small $\sigma$ (see Panel C), reflecting the increased probability that a product with a buy recommendation is controversial.  [^27] There is also a positive effect of higher $Q$, reflecting a change in the value of the outside option. The values of $\sigma$ and $\beta$ determine which effect dominates.
Corollary is an immediate implication of the following statement:
Lemma A.3.
Suppose Assumption 1 holds.
  • (i)
    If $03\sigma-\beta-\sigma^{2}+\sigma^{2}\beta>$ 0, then the value of the recommendation system is decreasing in $Q$. In particular, this is the case when $\sigma=1$.
  • (ii)
    If $03\sigma-\beta-\sigma^{2}+\sigma^{2}\beta<$ 0, then $Q^{*}:=\frac{3\sigma-\beta-\sigma^{2}+\sigma^{2}\beta}{4\sigma-4\beta-4\sigma^{2}+4\sigma^{2}\beta}\in(0,\frac{1}{2})$ maximizes $V(R)$.
Proof.
We first take the derivative of $V$ with respect to $Q$ to obtain
$\frac{\partial V}{\partial Q}=\frac{1}{\left(\sigma+1\right)^{2}}\left(\beta-3\sigma+4Q\sigma-4Q\beta+\sigma^{2}-4Q\sigma^{2}-\sigma^{2}\beta+4Q\sigma^{2}\beta\right)$
Setting this equal to zero and solving for the candidate solution we obtain
$Q^{\ast}=\frac{3\sigma-\beta-\sigma^{2}+\sigma^{2}\beta}{4\sigma-4\beta-4\sigma^{2}+4\sigma^{2}\beta}$
Next, the SOC reads
$4\sigma-4\beta-4\sigma^{2}+4\sigma^{2}\beta<0,$
where the l.h.s. of the SOC is the denominator of $Q^{\ast}$. Thus, a necessary condition for a positive interior solution is
$3\sigma-\beta-\sigma^{2}+\sigma^{2}\beta<0\text{.}$
This proves (i). To prove (ii), note that $03\sigma-\beta-\sigma^{2}+\sigma^{2}\beta<$ 0 implies $04\sigma-4\beta-4\sigma^{2}+4\sigma^{2}\beta<$ 0. Hence, the candidate solution is positive. It remains to show that $Q^{\ast}<\frac{1}{2}$. Recall that, by the requirements that the SOC holds and $Q^{\ast}>0$, we can focus on the case that both the numerator and the denominator in $Q^{\ast}$ are negative. Accordingly, the requirement that $\frac{3\sigma-\beta-\sigma^{2}+\sigma^{2}\beta}{4\sigma-4\beta-4\sigma^{2}+4\sigma^{2}\beta}<1/2$ becomes
$2\sigma-2\beta-2\sigma^{2}+2\sigma^{2}\beta<3\sigma-\beta-\sigma^{2}+\sigma^{2}\beta.$
Equivalently,
$\left(\sigma+1\right)\left(\sigma(1-\beta)+\beta\right)>0$
which always holds for $\beta\leq 1$.
A.3 Distinct Sender and Receiver Distributions
We prove the following result:
Proposition A.2.
Consider the model with distinct preference distributions of senders and receivers. Suppose the sender distribution $F$ satisfies Assumption 1. Then, all types accept recommendations. Further, $V(R)=\pi^{B}_{F}[\Delta_{O,F}^{B}-\Delta_{S,F}^{B}\mathbb{E}_{G}[i]]$ is
  1. (i)
    decreasing in $R$ if $\sigma<\frac{Q-(q_{1}-q_{2})\mathbb{E}_{G}[i]}{Q-(q_{2}-q_{1})\mathbb{E}_{G}[i]}$;
  2. (ii)
    increasing in $R$ if $\sigma>\frac{Q-(q_{1}-q_{2})\mathbb{E}_{G}[i]}{Q-(q_{2}-q_{1})\mathbb{E}_{G}[i]}$;
  3. (iii)
    constant in $R$ if $\sigma=\frac{Q-(q_{1}-q_{2})\mathbb{E}_{G}[i]}{Q-(q_{2}-q_{1})\mathbb{E}_{G}[i]}$.
Proof.
The value of the recommendation system can be written as
[…]
and the result from Proposition 3 translates accordingly so that we can rewrite $V(\mathcal{R})$ as
$\pi^{B}[\Delta_{O,F}^{B}-\Delta_{S,F}^{B}\mathbb{E}_{G}[i]]$ if |$\text{if }|\Delta_{S,F}^{B}|\leq 2\Delta_{O,F}^{B}$
$(1-\pi^{B})G(\tilde{i})\left[\Delta_{O}^{D}-\Delta_{S,F}^{D}\mathbb{E}_{G}[i\mid i\leq\tilde{i}]\right]+\pi^{B}(1-G(\tilde{i}))\left[\Delta_{O,F}^{B}-\Delta_{S,F}^{B}\mathbb{E}_{G}[i\mid i\geq\tilde{i}]\right]$ if $\text{if }\Delta_{S,F}^{B}<-2\Delta_{O,F}^{B}$
$\pi^{B}G(\tilde{i})\left[\Delta_{O,F}^{B}-\Delta_{S,F}^{B}\mathbb{E}_{G}[i\mid i\leq\tilde{i}]\right]+(1-\pi^{B})(1-G(\tilde{i}))\left[\Delta_{O}^{D}-\Delta_{S,F}^{D}\mathbb{E}_{G}[i\mid i\geq\tilde{i}]\right]$ if $\text{if }\Delta_{S,F}^{B}>2\Delta_{O,F}^{B}.$
Since $F$ satisfies Assumption 1, all receivers will accept the recommendation so that we are in the case $|\Delta_{S,F}^{B}|\leq 2\Delta_{O,F}^{B}$. Thus, we have
$V(\beta)$ $=\pi^{B}[\Delta_{O,F}^{B}-\Delta_{S,F}^{B}\mathbb{E}_{G}[i]]$
$=q_{H}+\beta Q-(q_{H}+2\beta Q)(q_{H}+Q)+(q_{2}-q_{1})(q_{H}-\beta(1-2Q))\mathbb{E}_{G}[i]$
and therefore
$V^{\prime}(\beta)=Q-2Q(q_{H}+Q)+(q_{2}-q_{1})(1-2Q)\mathbb{E}_{G}[i]$
Replacing $q_{H}=((1-2Q)\sigma)/(\sigma+1)$ and rearranging the inequalities $V^{\prime}(\beta)>0$ and $V^{\prime}(\beta)<0$ yields the statement in the proposition.
A.4 Multiple Recommendation Levels
Proof of Proposition 10
The posteriors are given as follows. Following a buy recommendation,
$p_{H}^{B}(R)$ $=\frac{q_{H}}{q_{H}+q_{1}(1-F(R_{2}-1/2))+q_{2}F(1/2-R_{2})},$
$p_{1}^{B}(R)$ $=\frac{q_{1}(1-F(R_{2}-1/2))}{q_{H}+q_{1}(1-F(R_{2}-1/2))+q_{2}F(1/2-R_{2})},$
$p_{2}^{B}(R)$ $=\frac{q_{2}F(1/2-R_{2})}{q_{H}+q_{1}(1-F(R_{2}-1/2))+q_{2}F(1/2-R_{2})}$
$p_{L}^{B}(R)$ $=0.$
A don’t-buy recommendation results in a posterior of
$p_{H}^{D}(R)$ $=0,$
$p_{1}^{D}(R)$ $=\frac{q_{1}F(R_{1}-1/2)}{q_{1}F(R_{1}-1/2)+q_{2}(1-F(1/2-R_{1}))+q_{L}},$
$p_{2}^{D}(R)$ $=\frac{q_{2}(1-F(1/2-R_{1}))}{q_{1}F(R_{1}-1/2)+q_{2}(1-F(1/2-R_{1}))+q_{L}}$
$p_{L}^{D}(R)$ $=\frac{q_{L}}{q_{1}F(R_{1}-1/2)+q_{2}(1-F(1/2-R_{1}))+q_{L}}.$
Finally, the recommendation neither to buy nor not to buy yields
$p_{H}^{N}(R)$ $=0$ $\quad p_{2}^{N}(R)$ $=\frac{q_{2}\Gamma_{2}}{q_{1}\Gamma_{1}+q_{2}\Gamma_{2}}$
$p_{1}^{N}(R)$ $=\frac{q_{1}\Gamma_{1}}{q_{1}\Gamma_{1}+q_{2}\Gamma_{2}}$ $\quad p_{L}^{N}(R)$ $=0.$
where $\Gamma_{1}\equiv F(R_{2}-1/2)-F(R_{1}-1/2)$ and $\Gamma_{2}\equiv F(1/2-R_{1})-F(1/2-R_{2})$. For the symmetric case of Assumption 1, setting $F(1/2-R_{2})=1-F(R_{2}-1/2)=\beta_{2}$ and $F(1/2-R_{1})=1-F(R_{1}-1/2)=\beta_{1}$, this reduces to
$p_{H}^{B}$ $=\frac{q_{H}}{q_{H}+2Q\beta_{2}},$ $\quad p_{H}^{N}=0,\quad\quad\quad p_{H}^{D}$ $=0,$
$p_{1}^{B}$ $=\frac{q_{1}\beta_{2}}{q_{H}+2Q\beta_{2}},$ $\quad p_{1}^{N}=\frac{q_{1}}{q_{1}+q_{2}},\quad\quad\quad p_{1}^{D}$ $=\frac{q_{1}(1-\beta_{1})}{1-q_{H}-2Q\beta_{1}},$
$p_{2}^{B}$ $=\frac{q_{2}\beta_{2}}{q_{H}+2Q\beta_{2}},$ $\quad p_{2}^{N}=\frac{q_{2}}{q_{1}+q_{2}},\quad\quad\quad p_{2}^{D}$ $=\frac{q_{2}(1-\beta_{1})}{1-q_{H}-2Q\beta_{1}},$
$p_{L}^{B}$ $=0,$ $\quad p_{L}^{N}=0,\quad\quad\quad p_{L}^{D}$ $=\frac{q_{L}}{1-q_{H}-2Q\beta_{1}}.$
Next, we determine the indifferent type following an intermediate recommendation. Type $i$ will buy the product with the intermediate recommendation whenever
$\frac{q_{1}}{q_{1}+q_{2}}(1/2+i)+\frac{q_{2}}{q_{1}+q_{2}}(1/2-i)\geq q_{H}+q_{1}(1/2+i)+q_{2}(1/2-i)$
$\Leftrightarrow$ $i\left(\frac{(q_{1}-q_{2})(q_{H}+q_{L})}{q_{1}+q_{2}}\right)\geq\frac{1}{2}(q_{H}-q_{L})$
If $q_{1}=q_{2}$, all types buy if $q_{L}\geq q_{H}$ and none buy otherwise. If $q_{1}>q_{2}$, all types
$i\geq\frac{1}{2}\frac{(q_{H}-q_{L})(q_{1}+q_{2})}{(q_{1}-q_{2})(q_{H}+q_{L})}$
buy the product if $r=N$ [^28] The argument for $q_{1}<q_{2}$ is analogous. Taken together, we can define
$\tilde{i}_{M}:=\begin{cases}1/2&q_{1}=q_{2},q_{L}<q_{H}\\ -1/2&q_{1}=q_{2},q_{L}\geq q_{H}\\ \frac{1}{2}\frac{(q_{H}-q_{L})(q_{1}+q_{2})}{(q_{1}-q_{2})(q_{H}+q_{L})}&q_{1}\neq q_{2}.\end{cases}$
The value of the recommendation system therefore reads
$V(\beta_{2},\beta_{1})=$ $(q_{H}+2Q\beta_{2})\left(\frac{q_{H}+Q\beta_{2}}{q_{H}+2Q\beta_{2}}-(q_{H}+Q)\right)$
$+2Q(\beta_{1}-\beta_{2})\int_{\tilde{i}_{M}}^{1/2}1/2(q_{L}-q_{H})+i\frac{(1-2Q)(q_{1}-q_{2})}{2Q}dF(i),$
where the first part corresponds to the value of accepting the highest recommendation and the second part comes from buying following an intermediate recommendation. Depending on parameters, $\tilde{i}_{M}$ changes so that some, all or no types buy following an intermediate recommendation. Taking the respective derivatives we obtain
$\frac{\partial{V}}{\partial{\beta_{2}}}$ $=Q(q_{L}-q_{H})F(\tilde{i}_{M})-(1-2Q)(q_{1}-q_{2})(1-F(\tilde{i}_{M}))\mathbb{E}[i\mid i\geq\tilde{i}_{M}]$
$\frac{\partial{V}}{\partial{\beta_{1}}}$ $=Q(q_{L}-q_{H})(1-F(\tilde{i}_{M}))+(1-2Q)(q_{1}-q_{2})(1-F(\tilde{i}_{M}))\mathbb{E}[i\mid i\geq\tilde{i}_{M}]$
Suppose $\sigma>1$ so that $q_{L}<q_{H}$. Then, $\frac{\partial{V}}{\partial{\beta_{2}}}<0$. Next, consider $\sigma<1$ so that $q_{L}>q_{H}$. Then, $\frac{\partial{V}}{\partial{\beta_{1}}}>0$.
Finally, suppose $q_{1}=q_{2}$. If $q_{L}<q_{H}$, then $\tilde{i}_{M}=1/2$, so that the derivatives read
$\frac{\partial{V}}{\partial{\beta_{2}}}$ $=Q(q_{L}-q_{H})<0$
$\frac{\partial{V}}{\partial{\beta_{1}}}$ $=0\text{.}$
For $q_{L}>q_{H}$, we obtain $\tilde{i}_{M}=-1/2$, so that the derivatives read
$\frac{\partial{V}}{\partial{\beta_{2}}}$ $=0$
$\frac{\partial{V}}{\partial{\beta_{1}}}$ $=Q(q_{L}-q_{H})>0.$
Thus, for $q_{L}<q_{H}$ any $\beta_{1}$ and thus any $R_{1}$ is optimal, whereas for $q_{L}>q_{H}$ any $\beta_{2}$ and thus any $R_{2}$ is optimal.
A.5 Multiple Recommendations
We now derive some auxiliary results for the case of multiple recommendations, before we prove Proposition 11.
After $b$ buy recommendations and $d$ don’t-buy recommendations, where $b+d>0$, the posterior reads
$p_{H}(b,d)$ =$=\begin{cases}\frac{q_{H}}{q_{H}+q_{1}(\phi_{1}(R))^{b}+q_{2}(\phi_{2}(R))^{b}}&\text{if }d=0\\ 0&\text{if }b=0\\ \end{cases}$
$p_{1}(b,d)$ $=\frac{q_{1}(\phi_{1}(R))^{b}(1-\phi_{1}(R))^{d}}{q_{1}(\phi_{1}(R))^{b}(1-\phi_{1}(R))^{d}+q_{2}(\phi_{2}(R))^{b}(1-\phi_{2}(R))^{d}}$
$p_{2}(b,d)$ $=\frac{q_{2}(\phi_{2}(R))^{b}(1-\phi_{2}(R))^{d}}{q_{1}(\phi_{1}(R))^{b}(1-\phi_{1}(R))^{d}+q_{2}(\phi_{2}(R))^{b}(1-\phi_{2}(R))^{d}}$
$p_{L}(b,d)$ =$=\begin{cases}\frac{q_{L}}{q_{1}(1-\phi_{1}(R))^{d}+q_{2}(1-\phi_{2}(R))^{d}+q_{L}}&\text{if }b=0\\ 0&\text{if }d=0\\ \end{cases}$
Lemma A.4.
Consider the model with multiple recommendations. Suppose Assumption 1 holds and consider any threshold $R\in(0,1)$.
  1. (i)
    If the receiver obtains only buy (don’t buy) recommendations, then, as the number of recommendations increases, her posterior belief converges to $p_{H}=1$ ($p_{L}=1$).
  2. (ii)
    If the receiver obtains mixed recommendations, then $p_{1}(b,d)/p_{2}(b,d)=q_{1}/q_{2}$ and $p_{H}(b,d)=p_{L}(b,d)=0$ for any $b,d>0$.
Proof.
(i) Suppose the receiver gets only buy recommendations and that $b\rightarrow\infty$. Then, $\phi_{1}(R)=\phi_{2}(R)\in(0,1)$ if $R\in(0,1)$, so that $\lim_{b\rightarrow\infty}(\phi_{i}(R))^{b}=0$ for $i=1,2$. Hence, $\lim_{b\rightarrow\infty}p_{H}(b,0)=1$. The argument for don’t-buy recommendations is analogous.
(ii) For any $b,d>0$, $p_{H}(b,d)=p_{L}(b,d)=0$ obviously holds. Further, because $\phi_{1}(R)=\phi_{2}(R)$ by Assumption 1,
$p_{i}(b,d)=\frac{q_{i}}{q_{1}+q_{2}}$
for $i=1,2$, yielding statement (ii).
Optimal Receiver Behavior with Infinite Learning
Lemma A.5.
Consider the “infinite learning” recommendation system with interior threshold $R\in(0,1)$. Then,
  1. (i)
    any receiver type $i\in[-1/2,1/2]$ buys the product if it is good;
  2. (ii)
    no receiver type $i\in[-1/2,1/2]$ buys the product if it is bad;
  3. (iii)
    if the product is controversial,
    1. (a)
      and $\lambda>1$, all receiver types $i\geq\tilde{i}_{\infty}$ buy the product;
    2. (b)
      and $\lambda<1$, all receiver types $i\leq\tilde{i}_{\infty}$ buy the product.
    3. (c)
      and $\lambda=1$, all receiver types $i\in[-1/2,1/2]$ buy the product if $\sigma\leq 1$ and none buy it if $\sigma>1$.
Proof.
First, Lemma A.4 implies that infinite learning fully reveals the good and bad products, whereas the posteriors of the controversial product versions are $p_{i}(b,d)=q_{i}/(q_{1}+q_{2})$ for $b,d>0$ Hence, the expected payoff of buying the recommended good with mixed recommendations reads
$\frac{q_{1}}{q_{1}+q_{2}}(1/2+i)+\frac{q_{2}}{q_{1}+q_{2}}(1/2-i)=\frac{1}{2}+i\frac{q_{1}-q_{2}}{q_{1}+q_{2}}.$
Therefore, receiver $i$ buys the product with mixed recommendations if and only if
$\frac{1}{2}+i\frac{q_{1}-q_{2}}{q_{1}+q_{2}}\geq V^{0}(i)=q_{H}+q_{1}(1/2+i)+q_{2}(1/2-i).$ (16)
The indifferent consumer satisfies
$\frac{1}{2}+\tilde{i}_{\infty}\frac{q_{1}-q_{2}}{q_{1}+q_{2}}=q_{H}+q_{1}(1/2+\tilde{i}_{\infty})+q_{2}(1/2-\tilde{i}_{\infty})\text{ and thus}$
$\tilde{i}_{\infty}=\frac{(q_{1}+q_{2})(q_{1}+q_{2}+2q_{H}-1)}{2(q_{1}-q_{2})(1-(q_{1}+q_{2}))}$
Using $q_{1}=\frac{2Q\lambda}{\lambda+1}$, $q_{2}=\frac{2Q}{\lambda+1}$ and $q_{H}=((1-2Q)\sigma)/(\sigma+1)$ gives
$\tilde{i}_{\infty}=\frac{(\sigma-1)}{2(1+\sigma)}\frac{\lambda+1}{\lambda-1}$
Inserting this into (16) gives the result for part (iii) when $\lambda\neq 1$. For the case $\lambda=1$, the condition to accept the recommendation simplifies to
$\frac{1}{2}\geq q_{H}+Q\Leftrightarrow 1\geq 2q_{H}+(1-q_{H}-q_{L})\Leftrightarrow q_{L}\geq q_{H}\Leftrightarrow\sigma\leq 1.$
Finally, parts (i) and (ii) follow trivially.
Proof of Proposition 11
  1. (i)
    If $\lambda=1$ the value of the infinite learning recommendation system is given by
    $V_{\infty}=q_{H}(1-q_{H}-Q)+2Q\max\left\{\frac{1}{2}-q_{H}-Q,0\right\}.$
    The first term corresponds to the value generated when the product is good, so that the sender always buys it. The second term captures the value generated from a controversial product if it is bought, where the purchasing decision depends on the parameters. Hence, $V_{\infty}$ coincides with the value of the single-recommendation system for the optimal $\beta$, depending on whether $q_{L}$ or $q_{H}$ is bigger.
  2. (ii)
    If $\lambda>1$, only $i\geq\tilde{i}_{\infty}$ buy the product by Lemma A.5. Hence, no type buys a controversial product in the infinite-learning recommendation system if $\tilde{i}_{\infty}\geq 1/2$, which is equivalent to $\sigma\geq\lambda$. This inequality implies $\sigma>1$. Hence, in that case all receivers behave in the same way in both the optimal single-recommendation system (Proposition 5) and the infinite-learning system, buying only objectively good products. Conversely, if $\tilde{i}_{\infty}\leq-1/2$ all types buy a controversial product with the infinite-learning recommendation system. This can be rewritten as $\sigma\leq 1/\lambda$, which implies $\sigma<1$, so that all receivers buy all products except those that are objectively bad and therefore behave in the same way in the optimal single-recommendation system and the infinite-learning system.
  3. (iii)
    The proof is analogous to case (ii) and thus omitted.
Appendix B Interactive Plots of $V(R)$ and $D(R)$
In what follows, we describe how to generate interactive plots of $V(R)$ and $D(R)$ under Assumption 2. The numerical analysis is based on a self-contained R script using the shiny and ggplot2 packages. The script allows users to manipulate model parameters in real time and visualize the corresponding plots through a simple web-based interface.
Requirements
  • R (version 4.0 or higher recommended)
  • RStudio (optional but recommended)
The script requires the R packages shiny and ggplot2, both of which are automatically checked for and installed if necessary. No manual installation is required. Beyond that, you only need to download the file recommendations_simulations.R . This single R script contains the entire code for both value and demand plots.
Instructions for Use
  1. 1.
    Open RStudio.
  2. 2.
    Open the file recommendations_simulations.R .
  3. 3.
    Run the script (e.g., select all lines and press Ctrl + Enter or click “Run App”).
  4. 4.
    A Shiny application will launch in your Viewer pane or in a web browser.
  5. 5.
    Use the sliders to manipulate the parameters $\sigma$, $a$ and $Q$.
  6. 6.
    Navigate between the two tabs:
    • V(R): Value of the recommendation system
    • D(R): Demand for the recommended product
Technical Details
  • The $\sigma$ and $a$ sliders operate on a logarithmic scale centered around $1$ to enable easy exploration.
  • The $R$ thresholds are plotted from $0.001$ to $0.999$.
  • The app will dynamically update plots based on the current input values.